### v2

Abstract
We consider the problem of preprocessing an $n\times n$ matrix $\mathbf{M}$, and supporting queries that, for any vector $v$, returns the matrix-vector product $\mathbf{M}v$. This problem has been extensively studied in both theory and practice: on one side, practitioners have developed algorithms that are highly efficient in practice, whereas on the other side, theoreticians have proven that the problem cannot be solved faster than naive multiplication in the worst-case. This lower bound holds even in the average-case, implying that existing average-case analyses cannot explain this gap between theory and practice. Therefore, we study the problem for structured matrices. We show that for $n\times n$ matrices of VC-dimension $d$, the matrix-vector multiplication problem can be solved with $\widetilde{O}(n^{2})$ preprocessing and $\widetilde{O}(n^{2-1/d})$ query time. Given the low constant VC-dimensions observed in most real-world data, our results posit an explanation for why the problem can be solved so much faster in practice. Moreover, our bounds hold even if the matrix does not have a low VC-dimension, but is obtained by (possibly adversarially) corrupting at most a subquadratic number of entries of any unknown low VC-dimension matrix.
In turn, our results also yield the first non-trivial upper bounds for many applications. In previous works, the online matrix-vector (OMv) hypothesis (conjecturing that quadratic time is needed per query, even over the boolean semi-ring) was used to prove many conditional lower bounds, showing that it is impossible to compute and maintain high-accuracy estimates for shortest paths, Laplacian solvers, effective resistance, and triangle detection in graphs subject to node insertions and deletions in subquadratic time. Yet, via a reduction to our matrix-vector-multiplication result, we show we can maintain the aforementioned problems efficiently if the input is structured, providing the first subquadratic upper bounds in the high-accuracy regime.
keywords:
Online Matrix-Vector Multiplication, Boolean Matrix Multiplication, VC dimension, Pollard Pseudodimension, Parameterized Complexity, Dynamic Laplacian Solver
footnotetext: Work done while at Princeton University.
1 Introduction
Computing sequential matrix-vector products is a workhorse of many iterative algorithms in machine learning. In optimization (Renegar, 1996; Vempala et al., 2020; Lin et al., 2023a; Chaudhari et al., 2024), computational geometry (Har-Peled, 2009; Welzl, 1992; Fisikopoulos and Penaranda, 2016), online algorithms (Lin et al., 2024; Anand and Qu, 2024; Murray et al., 2021), and dynamic algorithms (Liu, 2024; Jiang et al., 2023; Anand et al., 2024b; Jin and Xu, 2022; Lin et al., 2023b), sequential matrix-vector products are essential subroutines underlying performant algorithms. The current learning revolution is powered by hardware specifically designed to perform such products, as both neural network evaluation and back-propagation require repeated matrix-vector products (Sørensen, 2012; Rumelhart et al., 1986). Hence, matrix-vector products are arguably one of the most important and fundamental subroutines, with any complexity improvement having wide-ranging implications, and is thus a prevalent research topic in both theory and practice.
The problem can be modeled via the following data structure task: construct a data structure that preprocesses a given $n\times n$ matrix $\mathbf{M}$. After preprocessing, we want to multiply vectors with $\mathbf{M}$ faster than naive matrix-vector multiplication. Depending on the application, this matrix $\mathbf{M}$ is fixed or the data structure may need to handle changes to $\mathbf{M}$. For instance, when calculating probabilities of a random walk (e.g., in the PageRank algorithm (Page et al., 1999)), performing the power-method, and evaluating a neural network, the matrix is fixed. On the other hand, in convex optimization problems, such as in solving a linear or semi-definite program, applying Newton-Raphson’s method, and training a neural network, the matrix is either the inverse of a Hessian, or is given by the network’s weights, which changes from one iteration to the next.
Beyond speedups due to improving hardware accelerators, practitioners have made tremendous progress in accelerating the computation of matrix-vector products through heuristics that run in $\mathrm{nnz}(\mathbf{M})$ (the number of non-zero entries of $\mathbf{M}$) worst-case time, but which are much faster in practice. This has led to speedups in training graph neural networks (Alves et al., 2024) and other learning algorithms (Floros et al., 2024).
While practitioners have had success in accelerating the computation of matrix vector products, from a theory perspective there are substantial lower bounds. Gronlund and Larsen (2015) showed that any $\mathrm{poly}(n)$-space data structure for arbitrary matrix-vector multiplication over sufficiently large fields (such as $\mathbb{R}$) require $\Omega(n^{2}/\log n)$ time. In particular, any “generic” algorithm that does not make assumptions about the specific field needs $\Omega(n^{2}/\log n)$ time. This quadratic lower bound was also proven for arithmetic circuits (Frandsen et al., 2001). While these are worst-case lower bounds, they also hold for the average case (Henzinger et al., 2022). The only non-trivial upper bounds beating quadratic time are over the Boolean semi-ring (i.e., $\{0,1\}$ with $x\oplus y=\min(1,x+y)$). Here, the famous mailman algorithm (Liberty and Zucker, 2006) achieves $O(n^{2}/\log n)$ runtime. In a line of work that reduced the problem to finding neighborhoods in a graph and smaller algebraic products, this query complexity was improved to $O(n^{2}/\log^{2}n)$ (Williams, 2007) and $O(n^{2-o(1)})$ (Larsen and Williams, 2017; Chakraborty et al., 2018; Abboud et al., 2024). However, none of these algorithms are truly subquadratic, i.e., $O(n^{2-\epsilon})$ for some constant $\epsilon>0$.
Moreover, while the $\Omega(n^{2})$ lower bounds only hold for fields, even over the Boolean semi-ring it is conjectured that no truly subquadratic time algorithm exists (Henzinger et al., 2015), referred to as the OMv (online matrix-vector multiplication) conjecture. This has led to a line of works that use the conditional hardness of OMv to prove tight time lower bounds for a host of dynamic algorithms, such as dynamic matrix inversion (Brand et al., 2019a, b), dynamic subgraph connectivity (Henzinger and Neumann, 2016), dynamic regression (Jiang et al., 2023), dynamic range mode (Jin and Xu, 2022), Erickson’s problem and Langerman’s problem (Jin and Xu, 2022), as well as high-dimensional generalizations of Klee’s measure problem (Yildiz and Suri, 2012).
In summary, there is a substantial difference between the perspectives of this problem in theory and practice: there are heuristics that are fast in practice, but from a theoretical perspective the problem is hard, and an entire subarea of fine-grained-complexity has been built on this hardness. Moreover, average-case analysis cannot explain this difference between theory and practice, as even the average-case is provably hard. In this work, we resolve this conflict. Since the average-case (i.e., random non-structured input) is hard, the observed efficiency of practical algorithms must stem from some inherent structure in the real-world input data. Thus, we must study the complexity for structured inputs.
A popular measure for structural complexity is the Vapnik-Chervonenkis (VC) dimension, which finds many applications in machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019), structural graph theory (Nguyen et al., 2024; Alon et al., 2016; Karczmarz and Zheng, 2024), computational geometry (Har-Peled, 2009; B. Chazelle, 1989; Fisikopoulos and Penaranda, 2016) (see the preliminaries in Section 2 for a definition of the VC-dimension). Many structured objects studied by theoreticians have inherently low VC dimension. For instance, the hypothesis class of intervals on $\mathbb{R}$, star and interval graphs, and rank-one projection matrices, have VC-dimension $2$, planar graphs have VC dimension 4, and any $H$-minor free graph has VC dimension at most $|V(H)|-1$. Beyond being an established complexity measure in theory, it has also been empirically verified that real-world data has low VC dimension (Coudert et al., 2024). This observation can be explained via the following structural characterization of matrices with low VC dimension:
[…]
Consider a hereditary class of $0/1$-matrices $\mathcal{M}$, meaning that $\mathcal{M}$ is closed under row/column deletion (i.e., each matrix $\mathbf{M}\in\mathcal{M}$ is still in $\mathcal{M}$ after deleting a row or column). If $\mathcal{M}$ is non-trivial, (i.e., does not contain every possible matrix), then there exists an absolute constant $c\in\mathbb{N}$ such that the VC-dimension of any matrix in $\mathcal{M}$ is at most $c$.
Presumably, whatever unknown structure is common in real-world matrices should not be lost when deleting rows/columns. We note that section 1 does not preclude other matrices, that do not satisfy this hereditary property, from having constant VC dimension, e.g., adjacency matrices of minor-free graphs.
1.1 Our Results
While theoretical objects such as grid-graphs, intersection graphs, and kernel-matrices have a low VC-dimension, they are not perfect representations of real-world data. For instance, while city street networks in the US are predominantly grid-graphs, there are always exceptions in form of bridges or a few diagonal streets. Exceptions can also occur due to errors in measurements or other corruptions. To capture such almost low VC-dimension objects, we define the “corrupted VC-dimension.”
[…]
We say that a set system $\mathcal{F}$ on a ground set of size $n$ has “corrupted VC-dimension $d$” if there is another set system $\mathcal{F}^{\prime}$ of VC-dimension $\leq d$ such that $\mathcal{F}$ can be obtained from $\mathcal{F}^{\prime}$ by adding and/or removing (i.e., corrupting) at most $O(n^{1-1/d})$ elements to/from each set.
For matrices $\mathbf{M}\in\{0,1\}^{m\times n}$, the corrupted VC-dimension $d$ is equivalent to the existence of matrices $\mathbf{L}\in\{0,1\}^{m\times n}$ and $\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with $\mathbf{M}=\mathbf{L}+\mathbf{S}$ where $\mathbf{L}$ has VC-dimension $d$ and the corruption matrix $\mathbf{S}$ has at most $O(n^{1-1/d})$ non-zero entries per row, i.e., flipping at most $O(n^{1-1/d})$ bits per row of $\mathbf{L}$.
Note that merely the existence of such an $\mathcal{F}^{\prime}$ (i.e., $\mathbf{L}$) of VC-dimension $d$ suffices for $\mathcal{F}$ (i.e., $\mathbf{M}$) to have corrupted VC-dimension $d$. We do not need to know $\mathcal{F}^{\prime}$ (i.e., $\mathbf{L}$) or which sets have been corrupted (i.e., $\mathbf{S}$). Furthermore, set families of corrupted VC-dimension $d$ are strict generalizations of set families of VC-dimension exactly $d$. Therefore, while we state our main results for matrices with corrupted VC-dimension $d$, they also hold for general matrices with VC-dimension $d$.
We now state the theoretical guarantees of our accelerated matrix-vector multiplication algorithm in Theorem 1.1, where the runtimes are parameterized by the corrupted VC dimension measure.
Theorem 1.1.
(Static Online Matrix-Vector Multiplication). If a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then after an $\widetilde{O}(mn)$-time preprocessing, there is a data structure $\mathcal{D}$ that can compute $\mathbf{M}v$ for any $v\in\mathbb{R}^{n}$ in $\widetilde{O}(mn^{1-1/d}+n)$ time, with high probability.
The problem of approximating the VC-dimension of any set system within a $(2-\epsilon)$ accuracy for any $\epsilon>0$ is known to be $\Sigma_{3}^{P}$-hard (Mossel and Umans, 2002); however, for our theorem, the algorithm does not need to know or compute the VC-dimension to achieve this runtime.
This data structure is based on Gasieniec and Lingas (2003) which proposed an algorithm for Boolean matrix multiplication. Their algorithm computes the matrix-matrix product via $n$ matrix-vector products, thus implicitly providing a matrix-vector multiplication data structure as well. Their complexity depends on the weight of the minimum spanning tree defined with respect to the Hamming-distance between the rows of $\mathbf{M}$. While $O(n^{2})$ in worst-case, they conjecture this should be unlikely in real-world data. This data structure was recently independently rediscovered in the graph neural networks community (Alves et al., 2024) where they experimentally verified that the algorithm is indeed efficient on real-world inputs. Our Theorem 1.1 now gives a theoretical explanation as to why these practical algorithm behave so much more efficient in practice than the worst-case and average-case lower bounds of $\Omega(mn)$.
We obtain Theorem 1.1 via techniques from computational geometry. A line of works (Welzl, 1988; B. Chazelle, 1989; Matoušek, 1991; Har-Peled, 2009) consider the following geometric intersection data structure problem: given a set of points in $\mathbb{R}^{d}$, preprocess them. Then, given a convex polytope for each query, return whether any of the points intersect the polytope.
Matrix vector products can be interpreted as $O(m)$ intersection problems if we only care about Boolean outputs, because in the Boolean case $(\mathbf{M}v)_{i}\neq 0$ if and only if there is an intersection between the position of the non-zero elements in $v$ and the non-zero elements in the $i$-th row of $\mathbf{M}$. We show that the computational geometry techniques work even for non-boolean outputs, and we also extend them from bounded VC-dimension to our notion of corrupted VC-dimension. We provide details of this proof in Section 3.2.
At last, we remark that prior work for the geometric intersection problem had exponential preprocessing time (Welzl, 1992) or unspecified polynomial time (Matoušek, 1991).
In addition to our novel complexity bound, we also provide alternative algorithms to those presented in Gasieniec and Lingas (2003); Alves et al. (2024). We provide more details below.
Dynamic Setting
We extend Theorem 1.1 by allowing $\mathbf{M}$ to undergo row and column updates. The proof is given in Section A.2.
Theorem 1.2.
(Dynamic Online Matrix-Vector Multiplication). Given a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, there is a data structure $\mathcal{D}$ with $\widetilde{O}(mn)$ preprocessing time that supports row and column updates (insertions/deletions) to $\mathbf{M}$ in $\widetilde{O}(n)$ and $\widetilde{O}(m)$ time respectively. Upon querying the data structure $\mathcal{D}$ with a vector $v\in\mathbb{R}^{n}$, it outputs $\mathbf{M}v$ in $\widetilde{O}(mn^{1-1/d^{*}}+n)$ time, with high probability, where $d^{*}$ is the largest corrupted VC-dimension of $\mathbf{M}$ throughout the history of its updates.
In the special case where the updates to $\mathbf{M}$ keep its corrupted VC-dimension $d$ invariant, each matrix-vector product is computed with high probability in $\widetilde{O}(mn^{1-1/d}+n)$ time.
Extensions to Transposed Matrices.
Assume we have a small VC-dimension bound on $\mathbf{M}^{\top}$, then the VC-dimension of $\mathbf{M}$ could be much larger than that of $\mathbf{M}^{\top}$. If $d$ bounds the VC-dimension of $\mathbf{M}^{\top}$, then the best bound on the VC-dimension of $\mathbf{M}$ is $2^{d+1}$ (Assouad, 1983), so Theorem 1.1 would have only $\widetilde{O}(mn^{1-1/2^{d}})$ query time. To alleviate this substantial slow-down in Theorems 1.1 and 1.2, we extend the algorithms so their complexity is also bounded by the VC-dimension of $\mathbf{M}^{\top}$.
When $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $d^{\prime}$, we prove analogous versions of Theorems 1.1 and 1.2 that allow computing $\mathbf{M}v$ with high probability in $O(nm^{1-1/d^{\prime}}+m)$ time after a $\widetilde{O}(mn)$-time preprocessing. Thus, if $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$ and $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ has corrupted VC-dimension $d^{\prime}$, then after an $\widetilde{O}(mn)$-time preprocessing, we can compute $\mathbf{M}v$ in time $\min\{O(mn^{1-1/d}+n,nm^{1-1/d^{\prime}}+m)\}$. For this, we do not need to know $d$ or $d^{\prime}$, as the algorithm automatically achieves the minimum runtime.
Extension to Non-Boolean Matrices.
The Pollard pseudodimension is a popular extension of the VC-dimension to non-binary thresholds, which also finds many applications in theoretical machine learning (Shalev-Shwartz and Ben-David, 2014; Bartlett et al., 2019) and characterizations in structural graph theory (Karczmarz and Zheng, 2024) (see the preliminaries in Section 2 for a definition of the Pollard pseudodimension). For a real-valued matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, if $d$ bounds the Pollard pseudodimension of $\mathbf{M}$ and if $A$ bounds the number of thresholds of $\mathbf{M}$, then by lemma A.13, $\mathbf{M}v$ can be computed in time $\widetilde{O}(Amn^{1-1/d}+n)$. For $A\leq\widetilde{O}(1)$, this provides a subquadratic runtime.
1.2 Applications
A Scaling technique with Applications to Random Walks and PageRank.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be the adjacency matrix of an $n$-node graph and let $\mathbf{D}\in\mathbb{R}^{n\times n}$ be the diagonal matrix with vertex-degrees on the diagonal. Then the normalized Laplacian is $\widetilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$. Even though $\widetilde{\mathbf{L}}$ is not Boolean, we can still use Theorem 1.1 since $\widetilde{\mathbf{L}}v=\mathbf{D}^{-1/2}(w-\mathbf{A}w)$ for $w=\mathbf{D}^{-1/2}v$ and the product $\mathbf{A}w$ can be computed via Theorem 1.1. The same technique extends to random-walk probability matrices. In turn, this allows fast simulations of Markov chains to compute stationary distributions and iterative computation of the PageRank (Page et al., 1999) via Theorem 1.1.
Complexity-Theoretic Implications.
The Boolean matrix multiplication (BMM) conjecture states that no algorithm can multiply two boolean $n\times n$ matrices in $O(n^{3-\epsilon})$ time for constant $\epsilon>0$ (Basch et al., 1995; Williams and Williams, 2010). This conjecture is used to prove lower bounds on combinatorial algorithms (i.e., algorithms that do not make use of Strassen-style fast matrix multiplication methods), motivated by the property that fast matrix multiplication, while being theoretically fast (currently $O(n^{2.372})$ Williams et al. (2024)), is very slow in practice due to large constants hidden in $O$-notation. Hence BMM gives a lower bound on more practical algorithms and can be used to prove that any further complexity improvement would require fast matrix multiplication.
Theorem 1.1 implies that for matrices with corrupted VC-dimension $d$, BMM can be solved in $\widetilde{O}(n^{3-1/d})$ time, allowing for improved combinatorial upper bounds. The BMM problem was the original motivation for Gasieniec and Lingas (2003). Moreover, as demonstrated in Alves et al. (2024), this algorithm is also highly efficient in practice and does not encounter the large constant of fast matrix multiplication. For matrices of VC-dimension $1$, Theorem 1.1 achieves the optimal matrix-multiplication time of $\widetilde{O}(n^{2})$ (upto polylogarithmic factors). Through this view, Theorem 1.1 implies that problems with lower bounds based on BMM, may be solved more efficiently on structured inputs.
Corollary 1.3.
Suppose we are given Boolean matrices $\mathbf{M}_{1},\mathbf{M}_{2}\in\{0,1\}^{n\times n}$, such that $\mathbf{M}_{i}$ has corrupted VC-dimension $d_{i}$ and $\mathbf{M}_{i}^{\top}$ has corrupted VC dimension $d_{i}^{\prime}$ for $i\in\{1,2\}$. Let $s=\min\{d_{1},d_{2},d_{1}^{\prime},d_{2}^{\prime}\}$. Then, there is an algorithm for boolean matrix multiplication with in $O(n^{3-\frac{1}{s}})$ time complexity.
Another relevant conjecture is the OMv conjecture, which states that no algorithm can preprocess a Boolean matrix $M\in\{0,1\}^{n\times n}$ in polynomial time, and then answer $n$ queries that compute $\mathbf{M}v^{(i)}$ for an online sequence of Boolean vectors $v^{(1)},...,v^{(n)}\in\{0,1\}^{n}$ in total time $O(n^{3-\epsilon})$. Observe that here the vector $v^{(i+1)}$ is only given after the data structure returned $\mathbf{M}v^{(i)}$, hence the use of fast matrix multiplication to compute the matrix product $\mathbf{M}~{}[v^{(1)}|...|v^{(n)}]$ is ruled out.
This conjecture is used to prove conditional lower bounds for many dynamic problems, such as maintaining shortest paths, effective resistances, reachability, maximum flow, bipartite matching, triangle detection, and many others in graphs undergoing vertex insertions and deletions (Henzinger et al., 2015). Even with use of fast matrix multiplication, no algorithm can beat $O(n^{2})$ time on dense graphs, which for many of these problems is equivalent to re-solving them from scratch whenever the graph changes. Since Theorem 1.1 shows the OMv conjecture does not hold for structured matrices, it opens up improved dynamic algorithms.
We next list some examples of new dynamic algorithms obtained via this technique, whose upper bounds are only achievable due to their structured nature, otherwise violating either the BMM or OMv conjecture.
Implications for Dynamic Algorithms.
In the following, we state results for dynamic graph algorithms. These are data structures that maintain some graph property while the underlying graph changes over time. We defer the formal proofs for these results to Appendix B.
An important class of instances of linear equations that arise in practice has the form $\mathbf{L}x=b$, where $\mathbf{L}$ is the Laplacian of an undirected graph $G=(V,E)$ given by $\mathbf{L}\coloneqq\mathbf{D}-\mathbf{A}$ where $\mathbf{D}$ is a diagonal matrix such that $\mathbf{D}_{i,i}=\mathrm{deg}(i)$ for $i\in[n]$ where $|V|=n$, and $\mathbf{A}$ is the adjacency matrix of the graph. There are Laplacian solvers that run in time nearly linear in the number of edges of the underlying graph (Spielman and Teng, 2004). We study the problem of constructing a dynamic Laplacian solver for $\mathbf{L}_{t}x=b_{t}$ where the Laplacian changes over time through vertex updates to the graph, and each $b_{t}$ is an arbitrary new vector.
Previous upper bounds could only handle a $\operatorname{poly}(1/\epsilon)$ accuracy by maintaining spectral sparsifiers (Abraham et al., 2016; Bernstein et al., 2022) or vertex sparsifiers (Durfee et al., 2019; Chen et al., 2020). Moreover, exact or $O(\log 1/\epsilon)$-time dependencies are ruled out by the OMv conjecture (since no algorithm beats naive recomputation from scratch in $\widetilde{O}(|E|)=\widetilde{O}(n^{2})$ time via nearly-linear time Laplacian system solvers (Spielman and Teng, 2004)). We provide the first result faster than naive recomputation from scratch in the high-accuracy $\log(1/\epsilon)$ regime.
Theorem 1.4.
(Dynamic Laplacian Solver). There is a dynamic algorithm that, given a dynamic graph $G=(V,E)$ with corrupted VC-dimension bounded by $d$, maintains a Laplacian system solver. The data structure supports queries that receive a vector $b\in\mathbb{R}^{|V|}$ and error parameter $\epsilon>0$. Then, in $\widetilde{O}(n^{2-1/d}\log 1/\epsilon)$ time, the algorithm returns the (approximate) solution $x$ to $\mathbf{L}x^{*}=b$ where $\|x-x^{*}\|_{\mathbf{L}}\leq\epsilon\|x^{*}\|_{\mathbf{L}}$. Each vertex update to $G$ takes $\widetilde{O}(n)$ time.
Using the fast dynamic Laplacian solver, we maintain dynamic effective resistances of a graph, which is a critical computation for graph clustering (Alev et al., 2017), fault-tolerant computing (Ghosh et al., 2008), network analysis (Tizghadam and Leon-Garcia, 2008; Anand and Umans, 2023; Anand et al., 2024a), and biological systems (Klein et al., 2022), as it signifies the degree of connectivity between two vertices. The effective resistance $r_{G}(a,b)$ between vertex $a$ and vertex $b$ in a graph $G$ is a graph-theoretic analog of leverage scores that represents the energy needed to route one unit of electric flow from $a$ to $b$, and has a natural linear algebraic formulation: let $\mathbf{e}_{i}$ denote the $i$-th standard basis vector. Then, for all $a,b\in V$, $r_{G}(a,b)=(\mathbf{e}_{a}-\mathbf{e}_{b})^{\top}\mathbf{L}^{\dagger}(\mathbf{e}_{a}-\mathbf{e}_{b})$, where $\mathbf{L}$ is the Laplacian of $\mathbf{G}$. Finally, $\mathbf{L}^{\dagger}$ denotes the Moore-Penrose pseudoinverse of $\mathbf{L}$.
Theorem 1.5.
(Dynamic Effective Resistance). There is a dynamic algorithm that, given a dynamic graph $G=(V,E)$ with corrupted VC-dimension bounded by $d$, maintains effective resistances in $G$. The data structure supports queries that receive a pair of vertices $u,v\in V$ and error parameter $\epsilon>0$. Then, in $\widetilde{O}(n^{2-1/d}\log 1/\epsilon)$ time, the algorithm returns a $(1\pm\epsilon)$-approximation of the effective resistance. Moreover, each node update to $G$ in the dynamic data structure takes $\widetilde{O}(n)$ time.
Our dynamic matrix-vector multiplication data structure also leads to faster dynamic triangle detection algorithms for graphs with corrupted VC-dimension $d$, which are critical components of fraud detection systems and community-detection algorithms. No subquadratic time algorithm exists for vertex updates, conditional on BMM (Abboud and Vassilevska Williams, 2014), and also conditional on OMv (Henzinger et al., 2015). We show that these lower bounds can be broken on structured graphs.
Theorem 1.6.
(Dynamic Triangle Detection). There is a dynamic algorithm that, given a dynamic graph $G=(V,E)$ with corrupted VC-dimension $d$, maintains whether $G$ contains a triangle or not. Each vertex update to $G$ takes $O(n^{2-1/d})$ time and returns a Boolean indicator for whether $G$ contains a triangle or not.
Furthermore, our techniques lead to faster dynamic approximate algorithms for single-source distances. In this problem, we are tasked with finding the distances from a designated source vertex to every other vertex.
Theorem 1.7.
(Dynamic Approximate Single-Source Shortest Paths). There is a dynamic algorithm that maintains $(1+\epsilon)$-approximate single source distances on a dynamic unweighted undirected graph $G=(V,E)$. If the corrupted VC-dimension of $G$ is bounded by $d$, each vertex update to $G$ takes $\widetilde{O}(kn^{2-1/(2d)}/\epsilon)$ time, and querying the distances for any source node takes $O(n^{2-1/(2d)}/\epsilon)$ time.
Given a metric space $\mathcal{M}$ with $n$ points and a positive integer $k\leq n$, the $k$-center problem asks one to select $k$ points, referred to as centers, such that the maximum distance of any point in the metric space to its closest center is minimized. It is known that the $k$-center is NP-hard to approximate within a factor of $(2-\epsilon)$ for any $\epsilon>0$ (Hsu and Nemhauser, 1979). The best existing dynamic algorithms (Cruciani et al., 2024) only work with edge updates and use fast matrix multiplication as a black-box. We present an algorithm without fast matrix multiplication that supports more powerful vertex updates.
Theorem 1.8.
(Dynamic Approximate $k$-center). Given an unweighted undirected graph $G=(V,E)$, there is a dynamic algorithm for $(2+\epsilon)$-approximate $k$-center with node update time $O(kn^{2-1/2d}/\epsilon)$, where $d$ is a bound on the corrupted VC-dimension of $G$.
1.3 Further Related Work
Results for general Boolean matrix multiplication.
A line of work shows that the complexity of matrix-vector multiplication for general Boolean matrices (where the multiplication is over the Boolean semiring) can attain a subpolynomial speed-up. For instance, Williams (2007) showed that BMM can be computed in $O(n^{3}/\log n)$ time. Larsen and Williams (2017) extended this to $O(n^{3}/2^{\Omega(\sqrt{\log n})})$. This was further extended in Abboud et al. (2024) to $O(n^{3}/2^{\Omega(\sqrt[7]{\log n})})$.
Results for structured matrices.
When the underlying matrix is Vandermonde, Toeplitz, Hankel, or Cauchy, a body of work shows that they admit fast matrix-vector multiplications through convolutional transformations (Motzkin, 1951; Olshevsky and Shokrollahi, 2000; Pan and Tsigaridas, 2014). Orthogonal-vector (OV) matrices are structured matrices where each row and column receives a label $v\in\{0,1\}^{d}$ and $\mathbf{M}_{i,j}=1$ if and only if the corresponding two labels are orthogonal. Similar to our corrupted VC-dimension results, Alman and Vassilevska Williams (2020) show that such OV matrices allow for fast matrix-vector products if there are few corruptions to the matrix.
Another approach to handle structured inputs is through the algorithms with predictions regime, where some learning algorithm extracts the structure and can predict information about future vectors, leading to faster matrix-vector multiplication algorithms Henzinger et al. (2024); Brand et al. (2024).
Result for dynamic structured graphs.
Our dynamic algorithms hold for any structured graph, even if we do not know their structure. When the structure is known, there exist specialized dynamic algorithms tailored to the specific graph class. These include, for instance, interval graphs Crespelle (2019); Chen et al. (2024), geometric graphs defined from Kernels (Alman et al., 2020), planar graphs (Korhonen et al., 2024), dynamic planar graphs (Abboud and Dahlgaard, 2016), and H-minor free graphs (Dorn et al., 2012).
Matrix vector multiplication in optimization.
For convex optimization algorithms, there is a long history on developing data structures that facilitate fast matrix-vector products. Here the matrices are generally projection matrices that require computation of some matrix inverse. While historically, most research was on accelerating the maintenance of the matrix inverse (Karmarkar, 1984; Vaidya, 1989; Lee and Sidford, 2015; Cohen et al., 2021; Lee et al., 2019; Brand, 2020; Jiang et al., 2021, 2020a; Huang et al., 2022; Jiang et al., 2022; Lee et al., 2015; Jiang et al., 2020b), this research has progressed so far that the current bottleneck for the fastest linear program solvers are simple matrix-vector products (Brand et al., 2020, 2021).
2 Preliminaries
Notation.
Let $[n]=\{1,\dots,n\}$. We use $O(\cdot)$ to hide constant factors, and $\widetilde{O}$ to hide polylogarithmic factors. Let $\|\cdot\|_{1}$ denote the $\ell_{1}$ (Manhattan distance). For any matrix $\mathbf{M}$, let $\mathbf{M}_{i}$ and $\mathbf{M}_{:,i}$ denote the $i$’th column and $i$’th row (respectively) of $\mathbf{M}$. For a set $V$ we write $\mathbf{M}\in\mathbb{R}^{m\times V}$ for the $m\times|V|$ matrix where we can index columns by $x\in V$, i.e., $\mathbf{M}_{x}\in\mathbb{R}^{m}$. For any vector $x\in\mathbb{R}^{n}$, the Hamming weight of $x$ is the number of non-zero entries of $x$ also denoted by $\operatorname{nnz}(x)$. For Boolean vectors $x,y\in\{0,1\}^{n}$, the Hamming distance between $x$ and $y$ is given by $\|x-y\|_{1}$. Finally, for any vector $x\in\mathbb{R}^{n}$ and positive definite matrix $\mathbf{L}\in\mathbb{R}^{n\times n}$, define the $\mathbf{L}$-induced norm of $x$ by $\|x\|_{\mathbf{L}}=\sqrt{x^{\top}\mathbf{L}x}$.
VC-Dimension of Range Spaces.
A range space (also called a set system) is a pair $\mathfrak{R}=(X,R)$ where $X$ is a set and $R$ is a set of subsets of $X$. Let $\Pi_{R}(A)=\{A\cap r:r\in R\}$. The VC-dimension of $\mathfrak{R}$ is given by $\mathrm{VC}(\mathfrak{R})=\max\{|S|:S\subseteq X\text{ and }|\Pi_{R}(S)|=2^{|S|}\}$. The primal shatter function $\pi$ of $\mathfrak{R}$ is given by $\pi(m)=\max\{|\Pi_{R}(A)|:A\subseteq X,|A|=m\}\text{ for $m\geq 0$}$. Let $B$ be a set of ranges in $R$. Then, let $\Pi_{X}^{*}(B)=\{C:C\subseteq X\text{and $C$ is not crossed by any range in $Q$}\}$ and let $\Pi_{X}^{*}(B)$ be maximal. Then, for $m\geq 0$, the dual shatter function $\pi^{*}$ of the range space is given by $\pi^{*}(m)=\max\{|\Pi_{X}^{*}(B)|:B\subseteq R,|B|=m\}$.
VC-Dimension of Matrices.
Any matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ corresponds to a set family given by $\mathcal{F}_{\mathbf{M}}=\{\{j:\mathbf{M}_{i,j}=1\}:i\in[m]\}$, i.e., consider each row of $\mathbf{M}$ as the indicator vector of some set. Through it, we can encode $\mathbf{M}$ by the range space $\mathfrak{R}_{M}=([n],\mathcal{F}_{\mathbf{M}})$. Then, the VC-dimension of $\mathbf{M}$ is defined as the VC-dimension of $\mathfrak{R}_{\mathbf{M}}$. Here, a subset $S$ of the columns of $\mathbf{M}$ is said to be shattered if each of the $2^{|S|}$ many 0/1 strings appears in some row in the restriction of $\mathbf{M}$ to the columns in $S$. Then, the VC-dimension of $\mathbf{M}$ is the maximum size of a shattered subset of the columns of $\mathbf{M}$. Moreover, the VC-dimension $d_{m,n}$ of a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ satisfies $1\leq d_{m,n}\leq\log n$.
Pollard Pseudodimension of Matrices.
For a family of functions $\mathcal{F}$ mapping $V$ to $\{0,\dots,\ell-1\}$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the largest integer $d$ such that there exists $x_{1},\dots,x_{d}\in V$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any $d$-length bit vector $\vec{b}=(b_{1},\dots,b_{d})\in\{0,1\}^{d}$, there exists an $f\in\mathcal{F}$ such that for all $i\in[d],f(x_{i})\leq y_{i}$ if and only if $b_{i}=1$. For a real matrix $\mathbf{M}$, define the pseudo-dimension of $\mathbf{M}$, denoted by $\mathrm{Pdim}(\mathbf{M})$ by thinking of the rows of $\mathbf{M}$ as functions and taking the pseudo-dimension of the resulting class of functions. Specifically, if $\mathbf{M}$ is an $m\times n$ matrix, for each $i\in\{1,\dots,m\}$, define $f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}$ by $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$, and let $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\{f_{\mathbf{M},i}:i\in\{1,\dots,m\}\})$.
3 The Structural Complexity of Matrix-Vector Multiplication
We present a technical overview of the algorithms and the main theorems. We begin this section with a quick recap of previous work Björklund and Lingas (2001); Alves et al. (2024) who established a connection between matrix vector products and minimum spanning trees on $n$ points $\{0,1\}^{m}$, where the edge weights are given by the Hamming distance between two points (i.e., matrix $\mathbf{M}$ defines a collection of points).
Having established this connection, we prove one of our main results in Section 3.2. We prove via techniques from computational geometry that the weight of the minimum spanning tree must be bounded, if the matrix $\mathbf{M}$ has corrupted VC-dimension at most $d$. This then yields our main result Theorem 1.1.
3.1 Matrix Vector Products via Minimum Spanning Trees
We here recap the algorithm idea by Björklund and Lingas (2001), which was later independently rediscovered by Alves et al. (2024). We also explain how to extend the results from $\mathbf{M}$ to $\mathbf{M}^{\top}$.
Differential Compression.
Let $\mathbf{M}\in\{0,1\}^{m\times n}$ be a binary matrix, and let $\mathbf{M}_{i}$ denote the $i$’th column of $\mathbf{M}$ for $i\in[n]$. We differentially compress $\mathbf{M}$ by writing each column $\mathbf{M}_{x}$ as the sum of another column $\mathbf{M}_{y}$ and a change vector $\Delta_{x,y}$ such that $\mathbf{M}_{y}=\mathbf{M}_{x}+\Delta_{x,y}$ and $\Delta_{x,y}=\mathbf{M}_{x}-\mathbf{M}_{y}$.
If $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ are similar, then $\|\Delta_{x,y}\|_{1}$ likely to be smaller than the number of nonzero elements of $\mathbf{M}_{x}$. Therefore, it is more efficient to indirectly represent $\mathbf{M}_{x}$ with respect to $\mathbf{M}_{y}$, rather than through the matrix $\mathbf{M}$. This necessitates an algorithm that finds a suitable chain of $\Delta$’s to represent all columns of $\mathbf{M}$: for each column $\mathbf{M}_{x}$, identify a similar column $\mathbf{M}_{y}$ that characterizes the former, such that the Hamming weight of $\Delta_{x,y}$ required to represent $\mathbf{M}_{x}$ is minimized subject to $\mathbf{M}_{y}$.
Since the algorithm measures the Hamming distance for each pair of columns in $\mathbf{M}$, we can model this compression by an undirected weighted graph on $G$ on $n$ vertices, where vertex $i$ represents column $i$ of $\mathbf{M}$, and the weight of each edge $(x,y)$ corresponds to the Hamming distance between columns $x$ and $y$. We can then find a Minimum Spanning Tree (MST) of $G$, which by definition spans $G$ with the minimum sum of edge weights possible. Therefore, any MST of $G$ rooted at vertex $r$ defines a chain of $\Delta$’s that, starting at $r$, can represent all the columns of $\mathbf{M}$.
We formalize this intuition by defining the $\Delta$-labeled spanning tree of a matrix and its weight.
Definition 3.1 ($\Delta$-labeled spanning tree of matrix $\mathbf{M}$).
A $\Delta$-labeled tree is a tree $T=(V,E)$ with some explicit root $x$, where each edge $e\in E$ is directed away from $x$ and labeled by a vector $\Delta_{e}\in\mathbb{R}^{m}$. We say $T$ is a $\Delta$-labeled spanning tree for matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ if $\mathbf{M}_{x}=\vec{0}$ for root $x$, and for all edges $(i,j)\in E$ we have $\Delta_{i,j}\coloneqq\mathbf{M}_{j}-\mathbf{M}_{i}$.
Definition 3.2 (Weight of a $\Delta$-labeled spanning tree).
For any $\Delta$-labeled spanning tree of $\mathbf{M}$ denoted by $T$ with directed edge-set $E$, define $\mathrm{weight}(T)$ as the total number of non-zero entries given by $\sum_{e\in T}\operatorname{nnz}(\Delta_{e})$.
The minimum spanning tree of $\mathbf{M}$ is then the $\Delta$-labeled spanning tree with the smallest possible weight. Given any spanning tree of $\mathbf{M}$, we use the weight of the $\Delta$-labeled spanning tree to parameterize the runtime of the matrix-vector multiplication algorithm. Here, the lower the weight of the $\Delta$-labeled spanning tree, the faster the runtime.
Lemma 3.3 (Gasieniec and Lingas (2003); Alves et al. (2024)).
Given a $\Delta$-labeled spanning tree $T$ of $\mathbf{M}^{\top}\in\{0,1\}^{n\times m}$ and a vector $v\in\mathbb{R}^{n}$, there is an algorithm that computes $\mathbf{M}v$ in $O(\mathrm{weight}(T)+n+m)$ time.
We describe this procedure in Algorithm 1.
As a quick recap, the algorithm computes $\mathbf{M}v$ by starting at a root node $r\in T$ of the spanning tree and performing depth-first search (DFS). The algorithm initializes an output vector and computes $(\mathbf{M}v)_{r}$ in $O(n+m)$ time. For each edge $(x,y)$ traversed in the recursion tree, the algorithm computes $(\mathbf{M}v)_{y}=(\mathbf{M}v)_{x}+\Delta_{y,x}^{\top}v$. The complexity of this iterative procedure is proportional to $\mathrm{weight}(T)=\sum_{e\in E(T)}\delta_{e}$ where for each edge $e=(x,y)$ in the $\Delta$-labeled spanning tree, $\delta_{e}$ is the number of non-zeros in $\Delta_{e}$, i.e., the Hamming distance between row $x$ and row $y$ of $\mathbf{M}$. For completeness sake, we provide a proof for Lemma 3.3 in A.7.
Since the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}^{\top}$ can be different from the weight of the $\Delta$-labeled spanning tree of $\mathbf{M}$, we provide a new algorithm with an analogous theoretical guarantee in lemma 3.4.
Lemma 3.4.
Given a $\Delta$-labeled spanning tree $T$ of $\mathbf{M}\in\{0,1\}^{m\times n}$ and a vector $v\in\mathbb{R}^{n}$, there is an algorithm that can compute $\mathbf{M}v$ in $O(\mathrm{weight}(T)+n+m)$ time.
To utilize the $\Delta$-labeled spanning tree of $\mathbf{M}$, note that $\mathbf{M}v=\sum_{i=1}^{n}\mathbf{M}_{i}v_{i}$. For simplicity, consider the special case where $\mathbf{M}$ has only two columns, then $\mathbf{M}v=\mathbf{M}_{1}v_{1}+\mathbf{M}_{2}v_{2}=\mathbf{M}_{1}(v_{1}+v_{2})+\Delta_{1,2}v_{2}$. For more columns, this can be extended to require only a product with $\mathbf{M}_{1}$ and products with each $\Delta_{e}$ of the $\Delta$-labeled spanning tree. Thus the complexity is bounded by $O(m+n+\text{weight}(T))$.
We provide proofs for Lemma 3.4 in A.7.
Since we wish to minimize the time complexity, we compute a minimum spanning tree. In Alves et al. (2024) this was done naïvely, whereas in Björklund and Lingas (2001) this is done via Theorem 4.2 of Indyk and Motwani (1998).
Lemma 3.5 (Theorem 4.2 of Indyk and Motwani (1998)).
For any $\epsilon>0$, we can compute an $(1+\epsilon)$-approximate MST with respect to the Hamming distance of $n$ points in $\{0,1\}^{m}$ in $\widetilde{O}(mn^{1+1/(1+\epsilon)})$ time.
Theorem 3.6.
Given Boolean $m\times n$ matrix $\mathbf{M}$, we can construct in $\widetilde{O}(nm)$ time a $\Delta$-labeled spanning tree whose weight is at most a $O(\log n)$-factor larger than the minimum spanning tree (MST) of $\mathbf{M}$.
Proof 3.7.
[…]
3.2 Existence of Low Weight $\Delta$-labeled Minimum Spanning Trees
The Hamming distance between any two columns in $\mathbf{M}$ is at most $m$. Since there are $O(n)$ edges in the $\Delta$-labeled MST, this gives a coarse bound of $O(mn)$ which is achieved for Hadamard matrices. One of our main technical contributions is to prove that $\Delta$-labeled MST’s with low weight exist for structure matrices. We give a bound on the MST weight parameterized by the corrupted VC-dimension of the matrix.
Theorem 3.8.
If a matrix has corrupted VC-dimension $d$, then the weight of the corresponding MST is at most $O(mn^{1-1/d}\log n)$.
We show that for matrices of bounded VC-dimension, the minimum spanning tree $T$ has small weight. For this, we work directly with the definition of VC-dimension and use set terminology, which allows us to use relevant work from Welzl (1988); B. Chazelle (1989); Matoušek (1991); Har-Peled (2009). This previous work studied bounded VC-dimensions in the context of computational geometry. We here use their results to bound the complexity of matrix vector products.
Definition 3.9 (Set crossings).
Let $A$ and $r$ be sets. We say that $A$ crosses $r$ if the symmetric difference is non-empty, i.e. $A\triangle r\neq\emptyset$. In other words, $A$ crosses $r$ if neither $A\subseteq r$ nor $A\cap r=\emptyset$ holds.
Definition 3.10 (Spanning Tree).
Let $A\subset X$ be a finite set of elements of a range space $(X,R)$. A spanning tree on $A$ is an undirected (unrooted) tree with node set $A$.
Definition 3.11 (Crossing number $\kappa(T)$ of spanning tree $T$).
A crossing of a range $r\in R$ in $T$ is an edge in $T$ that is crossed by $r$ (i.e., an edge that has one endpoint in $r$ and the other endpoint in $[n]\setminus r$). The crossing number $\kappa_{r}(T)$ or a range $r\in R$ is the number of crossings of $r$ in $T$. The crossing number $\kappa(T)$ of $T$ is the maximum crossing number of a range $r\in R$ in $T$. In other words, $\kappa(T):=\max_{r\in R}\kappa_{r}(T)$.
This notion of a spanning tree corresponds to the notion of the $\Delta$-labeled spanning tree for matrix $\mathbf{M}$. The set system $(X,R)$ induced by $\mathbf{M}\in\{0,1\}^{m\times n}$ uses set $X=[n]$ (i.e., column indices) with the rows of $\mathbf{M}$ being the indicator vectors of $m$ subsets of $X$. Thus, a spanning tree $T_{X}$ for $(X,R)$ and a spanning tree $T_{M}$ for $\mathbf{M}$ both consist of edges that connect elements from $[n]$. Furthermore, the number of crossings for an edge $(x,y)\in T_{X}$ is exactly the number of indices $i$ where $(\mathbf{M}_{x})_{i}$ and $(\mathbf{M}_{y})_{i}$ differ. Thus the crossing number of a tree $T_{X}$ for $(X,R)$ is exactly the weight of the corresponding $\Delta$-labeled spanning tree for $\mathbf{M}$.
For the rest of this section, we will argue that range spaces $(X,R)$ of corrupted VC-dimension $d$ have spanning trees with small crossing number.
The literature on finding a spanning tree on $X$ with a low crossing number stems from a line of earlier works on computational geometry from the 1900s (Welzl, 1988; B. Chazelle, 1989; Matoušek, 1991; Har-Peled, 2009). They used the trees so preprocess a set of points in $\mathbb{R}^{d}$ to quickly answer queries where for any arbitrary convex polytope, the query returns a point within the polytope. Here, we transfer these techniques to the online matrix-vector multiplication problem and extend them to the notion of corrupted VC-dimension.
Lemma 3.12 (Lemma 4.1 of Welzl (1988)).
Let $(X,R)$ be a range space VC-dimension at most $d$. For every $A\subseteq X,|A|=n\geq 1$ and every multiset $Q$ of ranges in $R$ with $|Q|=m$, there are two elements $x,y\in A$ where $x\neq y$ such that the number of ranges in $Q$ crossing $(x,y)$ is at most $O(m\cdot n^{-1/d}\log n)$.
This lemma allows us to prove the existence of a $\Delta$-labeled spanning tree for matrices with bounded VC-dimension. We later extend this to corrupted VC-dimension.
Lemma 3.13.
Let $(X,R)$ be a range-space with $|X|=n,|R|=m$ of VC-dimension at most $d$. Then there exists a spanning tree $T$ of crossing number $\kappa(T)=O(mn^{1-1/d})$.
Proof 3.14.
Let $A_{0}=X$. Consider the following process where initially $i=0$. By lemma 3.12, there exists distinct $x_{i},y_{i}\in A_{i}$ such that the number of ranges in $R$ crossing $(x_{i},y_{i})$ is at most $O(m\cdot|A_{i}|^{-1/d}\log n)$. Letting $A_{i+1}\leftarrow A_{i}\setminus x_{i}$ and repeating, constructs a spanning tree $T$.
By lemma 3.12, the total number of ranges that cross the edges in $T$ is at most
$\sum_{i=1}^{n}O(m(n-i)^{-1/d}\log(n-i))$ $\leq O\left(m\log n\int_{1}^{n}x^{-1/d}\mathrm{d}x\right)$
$=O(mn^{1-1/d}\log n).$
which proves the claim.
We now extend Lemma 3.13 to corrupted VC-dimension. For this, we want to linearize the tree constructed in Lemma 3.13, i.e., we make sure the tree is a single line of vertices. This is to reduce the impact of corruptions, so that any single corruption does not affect the number of crossings too much.
Consider for example where $X=[n]$, $R$ is a multi-set of $m$ empty ranges (corresponding to an $m\times n$ zero matrix $\mathbf{M}$), and the spanning tree is a star with $1\in X$ at the center. The corruption matrix $\mathbf{S}$ now adds 1 to each of the empty sets in $R$ (equivalently, turning the first column of $\mathbf{M}$ to an all-1-column). This would still fit into the definition of corrupted VC-dimension since each set is corrupted by only one element. Now the number of crossings of each edge is $O(m)$ because $1$ is in each of the $m$ sets, but none of the other elements in $X$ are in any set, and there are $n-1$ edges each connecting $1$ to another element of $X$. Thus there are $O(m)$ crossings and the tree has $O(mn)$ crossings in total.
If, however, the tree had been a single line, then the corruption would affect at most 2 edges, and the number of crossings in the tree would go up by only $O(m)$.
Lemma 3.15.
For range space $(X,R)$, $|X|=n$ with spanning tree $T$, there exists a permutation $\pi\in S_{n}$ such that the crossing number of the corresponding spanning tree $T^{\prime}$ (which is just the line $\pi(1),\pi(2),...,\pi(n)$) is at most twice that of $T$.
Proof 3.16.
Enumerate all the ranges in $R$ by $R_{1},\dots,R_{m}$, where $R_{i}:=\{c:\mathbf{M}_{i,c}=1\}$ for $i\in[m]$.
The crossing number of $R_{i}$ in ${T}$ counts the number of edges in ${T}$ such that one endpoint of the edge is in $R_{i}$ and the other endpoint is in $[n]\setminus R_{i}$. Suppose the spanning tree $T$ on $X$ has a crossing number of $\kappa_{i}$ for just $R_{i}$ and let $\kappa=\sum_{i=1}^{m}\kappa_{i}$ be the crossing number of $T$ for $(X,R)$.
Perform a DFS traversal on ${T}$, and write the in-order traversal $\mathcal{I}$. We use the permutation $\pi(i)=\mathcal{I}_{i}$ for $i\in[n]$. Let $T^{\prime}$ be the corresponding linear tree.
We see then that this vertex path can (in the worst case) cross $R_{i}$ $2\kappa_{i}$ times, that is because DFS uses each edge $(x,y)$ of $T$ twice: once when visiting $y$ (going down one level in the tree) and once when it is done processing $y$ (DFS pops the current level from the stack).
Thus the total number of crossings $\kappa(T^{\prime})$ for the linear tree $T^{\prime}$ is $\kappa(T^{\prime})\leq\sum_{i=1}^{m}2\kappa_{i}\leq 2\kappa(T).$
Proof 3.17.
of Theorem 3.8   Suppose that $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$. Let $\mathbf{L}\in\{0,1\}^{m\times n},\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with $\mathbf{M}=\mathbf{L}+\mathbf{S}$ where $\mathbf{L}$ has VC-dimension $d$ and $\mathbf{S}$ has at most $O(n^{1-1/d})$ non-zero entries per row. These matrices exist by definition of corrupted VC-dimension, i.e., $\mathbf{S}$ describes the corruption.
Let $(X,R)$ be the set system induced by $\mathbf{L}$, i.e., $X=[n]$ and each row of $\mathbf{L}$ is the indicator vector of a set in $R$ with $|R|=m$. Then $(X,R)$ also has VC-dimension $d$ and by Lemma 3.13 there exists a spanning tree $T$ of $(X,R)$ with crossing number bounded by $O(mn^{1-1/d}\log n)$.
This also implies the existence of a spanning tree $T^{\prime}$ which is a single line that has crossing number $O(mn^{1-1/d}\log n)$ by Lemma 3.15.
Now consider the range space $(X,R_{c})$ for the corrupted matrix $\mathbf{M}$. The spanning tree $T^{\prime}$ has the same crossing number on $(X,R_{C})$ as on $(X,R)$, except for extra crossings induced by the corruption. We consider a range $r\in R_{c}$ and count the number of new crossings it has incurred due to the corruption: an edge $(x,y)$ only has a new crossing for some range $r\in R_{c}$ if $x$ or $y$ was corrupted in $r$. Since each $x\in X$ has at most 2 incident edges in $T^{\prime}$ (since it is a line) each corruption to $r$ leads to at most $2$ new crossings. Since each range has at most $O(n^{1-1/d})$ corruptions by definition of VC-corruption, the total number of new crossing over all ranges is at most $O(mn^{1-1/d}\log n)$.
Thus, $T^{\prime}$ is a spanning tree for $(X,R_{c})$ with crossing number bounded by $O(mn^{1-1/d}\log n)$. This spanning tree corresponds to a $\Delta$-labeled spanning tree for $\mathbf{M}$, as $X=[n]$ are the column indices and the number of crossings for an edge $(x,y)$ is exactly the number of entries where $\mathbf{M}_{x}$ and $\mathbf{M}_{y}$ differ.
Since the minimum spanning tree can only have lower weight, we get an upper bound of $O(mn^{1-1/d}\log n)$ on the weight of the MST.
We can now put everything together to prove Theorem 1.1. See 1.1
Proof 3.18.
Given matrix $\mathbf{M}$, preprocess it by constructing a $\Delta$-labeled spanning tree $T$ via Theorem 3.6. This takes $\widetilde{O}(mn)$ time.
When receiving a vector $v\in\mathbb{R}^{n}$, pass $v$ together with $T$ to the algorithm of Lemma 3.4. It returns the product $\mathbf{M}v$ in time $\widetilde{O}(m+n+\text{weight}(T))$.
By Theorem 3.8, if $\mathbf{M}$ has corrupted VC-dimension bounded by $d$, then the MST of $\mathbf{M}$ has weight bounded by $O(mn^{1-1/d})$. Since $T$ is a $O(\log n)$-approximation of the MST, answering a query takes $\widetilde{O}(mn^{1-1/d}+n)$ time.
3.3 Dynamic Matrices
At last, we outline how to handle matrices $\mathbf{M}$ that are allowed to update over time, i.e., Theorem 1.2. See 1.2
We sketch how to handle column insertions: Let $\mathbf{M}$ be the current matrix. We split it into smaller matrices $\mathbf{M}^{(0)},\mathbf{M}^{(1)},...,\mathbf{M}^{(\log n)}$, where each $\mathbf{M}^{(i)}$ contains at most $2^{i}$ columns of $\mathbf{M}$. When a new column is inserted to $\mathbf{M}$, we insert that column to $\mathbf{M}^{(0)}$. When any $\mathbf{M}^{(i)}$ contains more than $2^{i}$ columns, then all its columns are inserted to $\mathbf{M}^{(i+1)}$ and $\mathbf{M}^{(i)}$ is reset to an empty matrix. Thus, any $\mathbf{M}^{(i)}$ receives new columns at most once every $2^{i}$ iterations to $\mathbf{M}$.
When an $\mathbf{M}^{(i)}$ receives new columns, we initialize a new matrix-vector multiplication data structure (Theorem 1.1) on it. This takes $\widetilde{O}(m2^{i})$ time, but amortizes to $\widetilde{O}(m)$ time per insertion to $\mathbf{M}$ since it only happens at most once every $2^{i}$ insertions. This yields an amortized time of $\sum_{i=0}^{\log n}\widetilde{O}(m)=\widetilde{O}(m)$ time to maintain all $\mathbf{M}^{(i)}$ for $i=0,...,\log n$.
To answer a matrix-vector multiplication query $\mathbf{M}v$, we split $v$ into vectors $v^{(0)},...,v^{(\log n)}$ such that $\mathbf{M}v=\sum_{i=0}^{\log n}\mathbf{M}^{(i)}v^{(i)}$. For the time complexity, note that each $\mathbf{M}^{(i)}$ is a submatrix of $\mathbf{M}$. Thus, the the time complexity needed to multiply a vector by $\mathbf{M}^{(i)}$ is bounded by $\widetilde{O}(mn^{1-1/d})$.
For column deletions, we can ignore them until $\mathbf{M}^{(i)}$ is reinitialized: simply set $v_{j}=0$ for the columns that should have been deleted. We can handle row deletions updates to $\mathbf{M}$ in a similar way, by splitting the matrix again into $\log m$ matrices, each containing at most $2^{j}$ rows. The full proof is given in Section A.2.
4 Conclusion
In this work, we study the structural complexity of sequential matrix-vector multiplication. We propose a theoretical framework to study heuristic algorithms that have achieved tremendous practical success. Building on the VC-dimension literature, we propose the notion of a corrupted VC-dimension $d$ and provide structural characterizations in section 1. In theorem 1.1, we show that if a matrix $\mathbf{M}\in\{0,1\}^{m\times n}$ has corrupted VC-dimension $d$, then its matrix-vector product $\mathbf{M}v$ can be computed in $O(mn^{1-1/d})$ time, providing polynomial speed-ups over existing methods. Moreover, our algorithm maintains updates to $\mathbf{M}$ in $\widetilde{O}(m)$ and $\widetilde{O}(n)$ time. This leads to the first $O(n^{2-\epsilon})$-time algorithms for maintaining high-accuracy estimates for Laplacian solvers, effective resistance, and triangle detection in graphs with node insertions and deletions.
Our results motivate interesting future work: for instance, it would be interesting if alternate generalizations of the VC-dimension, such as the Natarajan dimension, can also parameterize the structural complexity of matrix-vector multiplication. Moreover, we conjecture that our algorithm can be used to provide combinatorial accelerations to the context-free grammar parsing and the $k$-clique detection problem in structured instances. Finally, for the non-Boolean setting where we provide subquadratic runtimes in matrices with bounded Pollard pseudodimension, we conjecture that the dependence of the number of thresholds $A$ in the runtime can be reduced to be polylogarithmic instead of linear.
5 Acknowledgements
This work was supported by NSF Grants DMS 2202961 and CCF 2338816. We also express our thanks to Ce Jin, Shunhua Jiang, Elia Gorokhovsky, and Jingtong Sun for insightful discussions.
Appendix A Online Matrix Vector Data Structure
In this section we prove all the intermediate lemmas required for our first two main results on subquadratic online matrix-vector multiplication. The proof that combined the technical lemmas was already given in Section 3.
See 1.1
See 1.2
A.1 Static Online Matrix Vector Multiplication Data Structure
See 3.1
See 3.2
See 3.3
Proof A.1.
Consider Algorithm 1 which takes as input $\mathbf{M}\in\{0,1\}^{m\times n}$, a $\Delta$-labeled spanning tree of $\mathbf{M}^{\top}$ (denoted by $T$), and a vector $v\in\mathbb{R}^{n}$.
We inductively prove that $\Phi=\mathbf{M}v$ over the number of visited vertices. This is true when only one vertex is visited, as that is the root $r$ and by definition $\Phi_{r}=(\mathbf{M}v)_{r}$.
For the inductive step, assume DFS visits the $(k+1)$st vertex $x$. Let $y$ be the parent of the vertex, then the algorithm assigns
$\Phi_{x}$ $=\Phi_{y}+\Delta^{\top}_{y,x}v$
$=(\mathbf{M}v)_{y}+(\mathbf{M}_{x}-\mathbf{M}_{y})v$
$=(\mathbf{M}v)_{u}.$
Here $\Phi_{y}$ is already computed since it was among the first $k$ visited vertices.
Runtime analysis. Let $\delta_{i,j}$ be the number of non-zero entries in $\Delta_{i,j}$ for all $(i,j)\in E(T)$. Initializing the vector $\Phi$ and populating its first element $\Phi_{r}$ takes $O(n+m)$ time. The cost of populating the remaining entries is proportional to the number of non-zero entries in each $\Delta_{x,y}$, since we can store each $\Delta$ via a sparser representation of only its non-zero entries and use this to compute $\Delta_{y,x}^{\top}v$ in time $O(\delta_{y,x})$. Thus, the total time complexity is $O(\sum_{(x,y)\in E(T)}\delta_{x,y})=O(\mathrm{weight}(T))$. Hence, the runtime of algorithm 1 is $O(\mathrm{weight}(T)+n+m)$, proving the lemma.
Algorithm 1 Online Matrix-vector multiplication (given a $\Delta$-labeled tree of $\mathbf{M}^{\top}$)
0:  Input matrix $\mathbf{M}\in\{0,1\}^{m\times n}$, and a $\Delta$-labeled tree of $\mathbf{M}^{\top}$ denoted by $T$ and rooted at $r$.
0:  Input vector $v\in\mathbb{R}^{n}$.
  Initialize: $\Phi=0^{m}$
  $\Phi_{r}=\mathbf{M}_{r}^{\top}v$
  Run DFS on $T$ from root $r$. Whenever DFS visits a vertex $x\in V$ via some edge $(y,x)$, thenset $\Phi_{x}=\Phi_{y}+\Delta_{x,y}v$
  Return $\Phi$
Definition A.2 (Path-$\Delta$ matrix $\mathbf{N}^{(T)}$).
Given a $\Delta$-labeled subtree $T=(V,E)$ with root $r$, define the path-$\Delta$ matrix $\mathbf{N}^{(T)}\in\mathbb{R}^{m\times V}$ such that for $u\in V$,
$\mathbf{N}_{u}^{(T)}\coloneqq\sum_{e\in P(r\to u)}\Delta_{e}.$
Here, $P(r\to u)$ denotes the edges of the unique (since $T$ is acyclic) directed path from $r$ to $u$.
Lemma A.3.
For $\Delta$-labeled tree $T=(V,E)$, and a vector $v\in\mathbb{R}^{V}$, let $\vec{\sigma}\in\mathbb{R}^{E}$ with $\vec{\sigma}_{(y,c)}=\sum_{z\in V(T_{c})}v_{z}$, where $T_{c}$ is the subtree rooted at $c\in V$. Let $\vec{\Delta}\in\mathbb{R}^{m\times E}$ be the matrix with columns being the labels of tree $T$. Then $\vec{\Delta}\vec{\sigma}=\mathbf{N}^{(T)}v$.
Proof A.4.
By the definition of $\mathbf{N}^{(T)}$, we have
$\mathbf{N}^{(T)}v$ $=\sum_{y\in V}\mathbf{N}^{(T)}_{y}\cdot v_{y}$
$=\sum_{y\in V}\sum_{e\in P(r\to y)}\Delta_{e}\cdot v_{y}$
Now consider how often any one edge $e$ shows up in this sum: an edge $e=(x,z)$ is in $P(r\to y)$ if and only if $z$ is an ancestor of $y$ (or $y=z$) (or, in other words, if $y$ is a descendant of $z$ (or $y=z$)). This condition can be written as $y\in T_{z}$. Thus we write the sum as
$\sum_{y\in V}\sum_{e\in P(r\to y)}\Delta_{e}\cdot v_{y}$ $=\sum_{(x,z)\in E}\sum_{y\in T_{z}}\Delta_{x,z}\cdot v_{y}$
$=\sum_{(x,z)\in E}\Delta_{x,z}\sum_{y\in T_{z}}v_{y}$
$=\sum_{(x,y)\in E}\Delta_{x,y}\vec{\sigma}_{y}$
$=\vec{\Delta}\vec{\sigma},$
which proves the claim.
Lemma A.5.
Let $T$ be a $\Delta$-labeled spanning tree of $\mathbf{M}$ where the root $r$ of $T$ has $\mathbf{M}_{r}=\vec{0}$, then $\mathbf{N}^{(T)}=\mathbf{M}$.
Proof A.6.
We prove this by induction over the width of $\mathbf{M}$.
Base Case.
For width 1, $T$ is just the root $r$, so $\mathbf{N}^{(T)}=\vec{0}=\mathbf{M}$.
Inductive Step.
Suppose the statement holds for matrices of width $<n$ and we now have a matrix $\mathbf{M}$ of width $n$. Let $L\subset V$ be the leaves of $T$. Let $T^{\prime}$ be the tree $T$ with these leafs removed, and let $\mathbf{M}^{\prime}\in\{0,1\}^{m\times(V\setminus L)}$ be matrix $\mathbf{M}$ with all columns corresponding to $L$ removed. Then for any $y\in V\setminus L$
$\mathbf{M}_{y}=\mathbf{M}^{\prime}_{y}\stackrel{{\scriptstyle(1)}}{{=}}\mathbf{N}^{(T^{\prime})}_{y}\stackrel{{\scriptstyle(2)}}{{=}}\mathbf{N}^{(T)}_{y}$
where equality (1) uses the induction hypothesis and equality (2) uses the fact that any path in the leaf-removed tree $T^{\prime}$ also exists in $T$.
For any $y\in L$ let $z$ be its ancestor in $T$. Therefore, $z\notin L$. Then, we have
$\mathbf{N}^{(T)}_{y}$ $=\mathbf{N}^{(T)}_{z}+\Delta_{z,y}$ (1)
$=\mathbf{N}^{(T^{\prime})}_{z}+\Delta_{z,y}$ (2)
$=\mathbf{M}^{\prime}_{z}+\Delta_{z,y}$ (3)
$=\mathbf{M}_{z}+(\mathbf{M}_{y}-\mathbf{M}_{z})=\mathbf{M}_{y},$
where equation 1 uses the fact that the path from root $r$ to $y$ is a path from root $r$ to $z$ with one extra edge $(z,y)$. Equation 2 uses the fact the path from $r$ to $z$ in $T$ also exists in $T^{\prime}$, and finally equation 3 follows by the induction hypothesis. Together, they prove the lemma.
See 3.4
Proof A.7.
Algorithm:
We compute $\mathbf{N}^{(T)}v$ via Lemma A.3 which requires us to compute the vector $\vec{\sigma}\in\mathbb{R}^{E}$ with $\vec{\sigma}_{(y,c)}=\sum_{z\in V(T_{c})}v_{z}$ for all $(y,c)\in E$ where $T_{c}$ is the subtree of $T$ rooted at $c$. This can be done in $O(n)$ time via a simple “dynamic programming on trees” algorithm: For any $c\in V$, let $C$ be its children. Then
$\vec{\sigma}_{(y,c)}$ $=\sum_{z\in V(T_{c})}v_{z}$
$=v_{c}+\sum_{x\in C}\sum_{z\in V(T_{x})}v_{z}$
$=v_{c}+\sum_{x\in C}\vec{\sigma}_{(c,x)}.$
So we can compute $\vec{\sigma}$ in $O(n)$ by starting at the leaf-edges, then propagating the sums up towards the root. At last, we multiply $\vec{\Delta}\vec{\sigma}$ in $O(\text{weight}(T))$ time as that is the number of non-zeros in $\vec{\Delta}$ (which is the matrix consisting of the tree labels as column-vectors).
Correctness.
Without loss of generality, assume that for root $r$ of tree $T$ we have $\mathbf{M}_{r}=\vec{0}$. Otherwise, append a $\vec{0}$-column to $\mathbf{M}$, and add edge $(n+1,r)$ to $T$ with label $\Delta_{n+1,r}=\mathbf{M}_{r}-\mathbf{M}_{n+1}=\mathbf{M}_{r}$, and make $n+1$ the new root of $T$. This increases the tree weight by at most $m$, resulting in an additive $O(m)$ in the time complexity.
Append another entry to $v$ so that the dimensions match and $\mathbf{M}v$ is well-defined. By lemma A.5, we thus have $\mathbf{N}^{(T)}v=\mathbf{M}v$.
A.2 Dynamic OMv Data Structure
In this section we give a black-box reduction from dynamic OMv to static OMv. In the static version, we are given a matrix $\mathbf{M}$ to preprocess. This matrix stays fixed, while we receive an online sequence of vectors and must repeatedly compute $\mathbf{M}v$. In the dynamic version, we can inform the data structure about changes to $\mathbf{M}$. In particular, we are allowed to add/remove/change rows and columns of the matrix.
Using the reduction from dynamic to static, we obtain Theorem 1.2.
See 1.2
The idea of this reduction is to split the columns of $\mathbf{M}$ into $\log n$ matrices $\mathbf{M}_{0},...,\mathbf{M}_{\log n}$ where $\mathbf{M}_{i}$ contains at most $2^{i}$ columns. Each $\mathbf{M}_{i}$ is used as input to a static OMv data structure which is reinitialized whenever $\mathbf{M}_{i}$ changes which occurs at most once every $2^{i}$ updates to $\mathbf{M}$. This allows to amortize the reinitialization cost over multiple updates.
Lemma A.8.
Assume there is a static OMv data structure for matrices of size $m\times n$ with preprocessing time $P(m,n)$ and query time $Q(m,n)$. Then there exists a dynamic OMv data structure supporting row and column insertions and deletions. The preprocessing time is $\widetilde{O}(P(m,n))$, the amortized row update time is
$O\left(\sum_{j=0}^{\log m}\frac{P(2^{j},n)}{2^{j}}\right)$
and for column updates
$O\left(\sum_{i=0}^{\log n}\frac{P(m,2^{i})}{2^{j}}\right)$
and query time is $\widetilde{O}(Q(m,n))$. This reduction runs several static OMv data structures, each receiving a submatrix of the dynamic input matrix.
We start by proving column updates only, then extend the reduction to support row updates as well.
Lemma A.9.
Assume there is a static OMv data structure for matrices of size $m\times n$ with preprocessing time $P(m,n)$ and query time $Q(m,n)$. Then there exists a dynamic OMv data structure supporting column insertions and deletions. The preprocessing time is $\widetilde{O}(P(m,n))$, the query time is $\widetilde{O}(Q(m,n))$, and the amortized update time is
$O\left(\sum_{i=0}^{\log m}\frac{P(m,2^{i})}{2^{i}}\right)$
This reduction runs several static OMv data structures, each receiving a submatrix of the dynamic input matrix.
Proof A.10.
Let $\mathbf{M}$ be the input matrix that changes over time.
Initialization.
We run $\log n$ instances of the static OMv data structure, where the $i$th instance is guaranteed to run on a matrix with at most $2^{i}$ columns. Let $\mathbf{M}^{(0)},...,\mathbf{M}^{(\log n)}$ be the respective input matrices to these instances. Initially, only the last ($i=\log n$) data structure receives the initial matrix $\mathbf{M}$ as input, and all other instances of the data structure initialize on an empty matrix. Hence initialization takes $O(P(m,n))$ time.
Insertions.
Whenever there is a column insertion to $\mathbf{M}^{(i)}$ we reinitialize the $i$th static OMv data structure on that new matrix. When $\mathbf{M}^{(i)}$ has more than $2^{i}$ columns, then all columns are removed (i.e., we set $\mathbf{M}^{(i)}=0$) and inserted to matrix $\mathbf{M}^{(i+1)}$. All column insertions to $\mathbf{M}$ are passed to $\mathbf{M}^{(0)}$. Observe that any $\mathbf{M}^{(i)}$ is updated every $2^{i}$ updates to $\mathbf{M}$, hence the amortized update time is
$O\left(\sum_{i=0}^{\log n}P(m,2^{i})/2^{i}\right).$
Queries.
We can answer queries by observing that
$\mathbf{M}=[\mathbf{M}^{(\log m)}|...|\mathbf{M}^{(1)}|\mathbf{M}^{(0)}].$
So, for a query vector $v$ we split it into corresponding pieces $v^{(0)},...v^{(\log m)}$ and query each $\mathbf{M}^{(i)}v^{(i)}$. The time for this is
$O(Q(m,n)\log n)$
because each $\mathbf{M}^{(i)}$ is a submatrix of $\mathbf{M}$ and we have $\log n$ such matrices.
Deletions.
When deleting a column of $\mathbf{M}$, we do not delete the column from the respective $\mathbf{M}^{(i)}$. Instead, all future query vectors $v$ will have an extra 0 entry for the deleted column.
When all columns from $\mathbf{M}^{(i)}$ are appended to $\mathbf{M}^{(i+1)}$ because of some insertions, then the deleted columns are not appended to $\mathbf{M}^{(i+1)}$. Hence whenever a static OMv data structure is initialized, it is a submatrix of the input matrix $\mathbf{M}$.
When there have been $22^{i}/$ 2 postponed deletions to any $\mathbf{M}^{(i)}$, then the columns are actually deleted and the respective static OMv data structure is reinitialized on the new $\mathbf{M}^{(i)}$. The remaining columns in $\mathbf{M}^{(i)}$ all exist in the current dynamic input $\mathbf{M}$, so again, the static OMv data structure receives a submatrix of $\mathbf{M}$ as input. For any one $i$, this takes amortized $O(P(n,2^{i})/2^{i})$ time which is subsumed by the complexity of handling insertions.
Observe that internally, our matrices $\mathbf{M}^{(i)}$ are always at most twice their intended size because of the yet to be removed columns. Thus the query complexity increases by at most a constant factor.
Proof A.11.
of Lemma A.8 (Row & Column Updates)   Here, we extend Lemma A.9 to support row updates as well. The reduction follows the same idea as Lemma A.9.
Let $\mathbf{N}$ be the dynamic input matrix. We run $\log n$ copies of the column update data structure Lemma A.9 on matrices $\mathbf{N}^{(0)},...,\mathbf{N}^{(\log n)}$, with the $i$th matrix having at most $2^{i}$ rows. When inserting new rows to $\mathbf{N}$, they are passed to $\mathbf{N}^{(0)}$. When $\mathbf{N}^{(i)}$ has more than $2^{i}$ rows, all rows are appended to $\mathbf{N}^{(i+1)}$. When rows are appended to any $\mathbf{N}^{(i)}$, the respective data structure is reinitialized. To handle deletions, we remove the respective row from the output until $22^{i}/$ 2 rows have been deleted from $\mathbf{N}^{(i)}$, then the rows are removed from $\mathbf{N}^{(i)}$ and the respective data structure is reinitialized.
The amortized row update time thus becomes
[…]
For column updates, each of the $\mathbf{N}^{(i)}$ receives a new column, so each of the $\log m$ column update data structures is updated which takes:
$O\left(\sum_{j=0}^{\log m}\sum_{i=0}^{\log n}P(2^{j},2^{i})/2^{i}\right)=O\left(\sum_{i=0}^{\log n}P(m,2^{i})/2^{i}\right)$
because geometric sums are bounded by their largest term in $O$-notation.
Finally, the query time is
$\widetilde{O}\left(\sum_{j=0}^{\log m}Q(2^{j},n)\right)=\widetilde{O}(Q(m,n)),$
which proves the lemma.
Proof A.12.
of Theorem 1.2 (Dynamic OMv for corrupted VC-dimension $d$)   The initialization cost of static OMv is $P(m,n)=\widetilde{O}(mn)$, and the query complexity is $Q(m,n)=\widetilde{O}(mn^{1-1/d})$ by Theorem 1.1. This complexity also holds for submatrices, i.e., $Q(a,b)=\widetilde{O}(mn^{2-1/d}+n)$ for any $a\times b$ sized submatrix.
Thus we have update time for column updates:
$O\left(\sum_{i=0}^{\log n}P(m,2^{i})/2^{i}\right)=\widetilde{O}\left(\sum_{i=0}^{\log n}m2^{i}/2^{i}\right)=\widetilde{O}(m)$
and for row updates we have
$O\left(\sum_{j=0}^{\log m}P(2^{j},n)/2^{j}\right)=\widetilde{O}\left(\sum_{j=0}^{\log m}2^{j}n/2^{j}\right)=\widetilde{O}(n)$
and for query time we have
$\widetilde{O}(Q(m,n))\leq\widetilde{O}(mn^{1-1/d}+n),$
which proves the theorem.
A.3 Extension to Structured Non-Boolean Matrices
In this section, we prove the following lemma.
Lemma A.13.
Suppose that the Pollard pseudodimension of the matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ is $d$, and that the number of thresholds is $A$. Then, after an $\widetilde{O}(Amn)$ time preprocessing, there is a data structure that, upon receiving a vector $v\in\mathbb{R}^{n}$, can return $\mathbf{M}v$ in time $O(Amn^{1-1/d}+An)$.
Recall the definition of the Pollard pseudodimension:
Definition A.14 (Pollard Pseudodimension).
For a family of functions $F$ mapping $V$ to $\{0,\dots,\ell-1\}$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the largest integer $d$ such that there exists $x_{1},\dots,x_{d}\in V$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any $d$-length bit vector $\vec{b}=(b_{1},\dots,b_{d})\in\{0,1\}^{d}$, there exists an $f\in\mathcal{F}$ such that $\forall i,f(x_{i})\leq y_{i}\iff b_{i}=1$.
As an example, consider an integer matrix and split each row into the different thresholds.
$\begin{bmatrix}1&2&4&2\end{bmatrix}\to\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 1&1&0&1\\ 1&1&1&1\end{bmatrix}$
Then, the right matrix has VC dimension $d$, if and only if the left matrix has Pollard pseudodimension $d$.
Definition A.15 (Multiball vectors).
Let $v\in V$ be a vertex of the graph $G$. Let $r$ be a real number and $\Delta$ be a set of distances $-\infty=\delta_{0}<\delta_{1}<\dots<\delta_{\ell-1}<\delta_{\ell}=\infty$. Define the multiball vectors as follows:
$\vec{\mathrm{MB}}(v,r,\Delta)=(y_{u})_{u\in V}\text{ where }y_{u}\in[\ell]\text{ is the smallest integer such that }u\in\vec{B}(v,r+\delta_{yu}).$
Let $\vec{\mathrm{MB}}_{G,\Delta}=\{\vec{MB}(v,r,\Delta)|v\in V,r\in\mathbb{R}\}$ be the set of all multiball vectors in $G$ with shifts $\Delta$.
Theorem A.16 (Karczmarz and Zheng (2024)).
For a $K_{h}$-minor-free graph $G$, for any set $\Delta$ of $\ell$ distances, the set of multiball vectors $\vec{\mathrm{MB}}_{G,\Delta}$ has pseudodimension at most $h-1$.
Lemma A.17 (Generalized Sauer-Shelah lemma).
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$. If the pseudodimension of $\mathcal{F}$ is at most $d$, then $|\mathcal{F}|\leq O(n^{d}\ell^{d})$.
Lemma A.18.
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$. Let $\mathcal{F}_{s}=\mathcal{F}\cap[\ell]^{S}$ be a restriction of $\mathcal{F}$ to $S$. Then, $\mathrm{Pdim}(\mathcal{F}_{S})\leq\mathrm{Pdim}(\mathcal{F})$.
Corollary A.19.
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$, and let $S$ be any subset of the ground set of $\mathcal{F}$. Suppose $\mathrm{Pdim}(\mathcal{F})=d$. Then, $|\mathcal{F}_{S}|=O(|S|^{d}\ell^{d})$.
Definition A.20.
For a real matrix $\mathbf{M}$, define the pseudo-dimension of $\mathbf{M}$, denoted by $\mathrm{Pdim}(\mathbf{M})$ by thinking of the rows of $\mathbf{M}$ as functions and taking the pseudo-dimension of the resulting class of functions. Specifically, if $\mathbf{M}$ is an $m\times n$ matrix, for each $i\in\{1,\dots,m\}$, define $f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}$ by $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$ and let $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\{f_{\mathbf{M},i}:i\in\{1,\dots,m\}\})$.
Lemma A.21.
Given any matrix $\mathbf{M}\in\mathbb{Z}^{n\times n}$, construct the Boolean matrix $\mathbf{N}\in\{0,1\}^{n\times n}$ such that $\mathbf{N}_{i,j}=\mathbbm{1}\{\mathbf{M}_{i,j}\geq 0\}$ for $i,j\in[n]$. Then, $\mathrm{VC}(\mathbf{N})\leq\mathrm{Pdim}(\mathbf{M})$.
Proof A.22.
Given a matrix $\mathbf{M}$, we have threshold functions $f_{\mathbf{M},i}$ such that $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$ and can use this this to define a set family
=$\mathcal{F}=\{f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}|i=1,\dots,m\},$
and $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\mathcal{F})$.
By definition, $\mathrm{Pdim}(\mathbf{M})$ is the largest integer $d$ such that $\exists x_{1},\dots,x_{d}\in\{0,\dots,\ell-1\}$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any vector $b_{1},\dots,b_{d}\in\{-1,1\}^{d}$, there is an $f\in\mathcal{F}$ such that $\forall i,\mathrm{sign}(f(x_{i})-y_{i})=b_{i}$.
Define the following $0/1$ profile of matrix $\mathbf{M}$. For each row $i$ of $\mathbf{M}$, define the binary function
$g_{\mathbf{M}.i}:\{1,\dots,n\}\to\{0,1\}$
by
$g_{\mathbf{M},i}(j)=\mathbbm{1}\{f_{\mathbf{M},i}(j)\geq 0\}\coloneqq\mathbbm{1}\{\mathbf{M}_{i,j}\geq 0\},$
and use this to define the set family
$\mathcal{G}=\{g_{\mathbf{M},i}:i=1,\dots,m\}.$
We want to show that if $\mathrm{Pdim}(\mathbf{M})$ is low ($\mathrm{Pdim}(\mathbf{M})=d$), then the VC-dimension of $\mathcal{G}$ is at most $d$. So, by way of contradiction, suppose $\mathrm{Pdim}(\mathcal{F})=d$ and $\mathrm{VC}(\mathcal{G})\geq d+1$.
Then, there must exist $S\subset\{1,\dots,n\}$ with $|S|=d+1$ that is shattered by $\mathcal{G}$. So, for every $0/1$ labeling of $c:S\to\{0,1\}$, there is some row $i$ for which $g_{\mathbf{M},i}(j)=c(j),\forall j\in\mathcal{S}$.
Also, $\exists x_{1},\dots,x_{d}\in\{0,\dots,\ell-1\}$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any vector $b_{1},\dots,b_{d}\in\{-1,1\}^{d}$, there is an $f\in\mathcal{F}$ such that $\forall i\in[d],\mathrm{sign}(f(x_{i})-y_{i})=b_{i}$.
Also, note that the condition
$g_{\mathbf{M},i}(j)=\mathbbm{1}\{f_{\mathbf{M},i}(j)\geq 0\}=c(j)$
is equivalent to
$\mathrm{sign}(f_{\mathbf{M},i}(j)-0)=b^{*}_{j},\text{ for }b^{*}_{j}=\begin{cases}+1,&c(j)=1,\\ -1,&c(j)=0\end{cases}$
but this just means that $S$ is pseudo-shattered by the class $\mathcal{F}$, where the thresholds $y_{1},\dots,y_{d}=0$ for all $j\in S$. So, this implies $\mathrm{Pdim}(\mathbf{M})=d+1$, which is a contradiction. So, $\mathrm{VC}(\mathcal{G})\leq\mathrm{Pdim}(\mathcal{F})$.
Lemma A.23.
Suppose that the Pollard pseudodimension of the matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ is $d$, and that the number of thresholds is $A$. Then, after an $\widetilde{O}(Amn)$ time preprocessing, there is a data structure that, upon receiving a vector $v\in\mathbb{R}^{n}$, can return $\mathbf{M}v$ in time $O(Amn^{1-1/d}+An)$.
Proof A.24.
Given the matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ with $A$ thresholds $\tau_{1},\tau_{2},\dots,\tau_{A}\in\mathbb{R}^{A}$, construct the Boolean matrix $\mathbf{N}\in\{0,1\}^{Am\times n}$ as follows: for row $i\in[n]$ of $\mathbf{M}$, define the block $\mathcal{B}^{(i)}\coloneqq\mathcal{B}(\mathbf{M}_{i})\in\{0,1\}^{A\times n}$ by
$\mathcal{B}(\mathbf{M}_{i})=\begin{bmatrix}\mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{1}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{1}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{1}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{1}\}}\\ \mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{2}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{2}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{2}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{2}\}}\\ &&\ddots&&\\ \mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{A}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{A}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{A}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{A}\}}\end{bmatrix}$
Then, matrix $\mathbf{N}\in\{0,1\}^{Am\times n}$ is given by
$\mathbf{N}=\begin{bmatrix}\mathcal{B}^{(1)}\\ \mathcal{B}^{(2)}\\ \vdots\\ \mathcal{B}^{(m)}\end{bmatrix}$
By lemma A.21, the VC-dimension of $\mathbf{N}$ is at most the Pollard pseudodimension of $\mathbf{M}$, given by $d$. Constructing $\mathbf{N}$ takes $O(mnA)$ time. Then, using theorem 1.1, we can preprocess $\mathbf{N}$ in $\widetilde{O}(Amn)$ time to support queries that return $\mathbf{N}v$ for any vector $v\in\mathbb{R}^{n}$ in time $\widetilde{O}(Amn^{1-1/d}+n)$. To convert $\mathbf{N}v\in\mathbb{R}^{Am}$ to $\mathbf{M}v\in\mathbb{R}^{m}$, iterate over each of the $m$ sets of $A$ entries of $\mathbf{N}v$. For $i\in[m]$, denote the $i$’th set of $A$ entries by $x^{(1)}_{i},x^{(2)}_{i},\dots,x_{i}^{(A)}$, where $x_{i}^{(j)}\coloneqq(\mathbf{M}v)_{A(i-1)+j}$. Then, defining $y_{i}$ as the linear combination of the threshold maps with the outputs $x_{i}^{(j)}$, we have $y_{i}\coloneqq\sum_{j=1}^{A}x_{i}^{(A)}\tau_{i}$. Therefore, $\mathbf{N}v$ is given by
$\mathbf{N}v=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{m}\end{bmatrix},$
which proves the claim.
Appendix B Applications
In this section, we consider some applications of our OMv result in Theorem 1.2. These applications have quadratic $\Omega(n^{2})$-time lower bounds, conditional on the OMv conjecture. We show that on structured graph this lower bound can be beaten.
B.1 High-accuracy Dynamic Laplacian Solver
See 1.4
We consider the setting where the Laplacian $\mathbf{L}$ corresponds to a graph with bounded VC-dimension. First, recall the notion of spectral sparsifiers.
Definition B.1 ($(1+\epsilon)$-spectral sparsifiers).
Let $\mathbf{L}_{X}$ denote the Laplacian of any undirected graph $X$. Then, a $(1\pm\epsilon)$-spectral sparsifier $H$ of a graph $G$ is a subgraph of $G$ such that for every vector $x\in\mathbb{R}^{n}$,
$(1-\epsilon)x^{\top}\mathbf{L}_{H}x\leq x^{\top}\mathbf{L}_{G}x\leq(1+\epsilon)x^{\top}\mathbf{L}_{H}x,$
Then, a result from Abraham et al. (2016) provides a construction for a dynamic spectral sparsifier under edge deletions and insertions with polylogarithmic amortized update time:
Theorem B.2 (Theorem 4.1 of Abraham et al. (2016)).
There exists a fully dynamic randomized algorithm with polylogarithmic update time for maintaining a $(1\pm\epsilon)$-spectral sparsifier $H$ of a graph $G$, with probability at least $1-1/n^{c}$ for any $0<\epsilon\leq 1$ and $c\geq 1$. Specifically, the amortized update time of the algorithm is $O(c\epsilon^{-2}\log^{3}\rho\log^{6}n)$ and the size of $H$ is $O(cn\epsilon^{-2}\log^{3}\rho\log^{5}n\log W+m\rho^{-1})$, where $1\leq\rho\leq m$ is a parameter of choice. Here, $W$ is the ratio between the largest and the smallest edge weight in $G$.
We use the following result for solving Laplacian systems in the static setting.
Lemma B.3 (Kyng and Sachdeva (2016)).
There is a randomized procedure that given any $n$-vertex $m$-edge graph $G$ with Laplacian matrix $\mathbf{L}_{G}$, and vector $b\in\mathbb{R}^{V}$ such that there exists an $x\in\mathbb{R}^{V}$ with $\mathbf{L}_{G}x=b$ computes ${\overline{x}}\in\mathbb{R}^{V}$ with $\|{\overline{x}}-x\|_{\mathbf{L}_{G}}\leq\epsilon\|x\|_{\mathbf{L}_{G}}$ in $\widetilde{O}(m\log\epsilon^{-1})$ with high probability.
Lemma B.4.
Let $G$ be a dynamic graph and let $\mathbf{A}$ be its dynamic adjacency matrix. Assume there is a dynamic OMv data structure for this $\mathbf{A}$ with update time $U(n)$ and query time $Q(n)$.
Then there exists a dynamic Laplacian System solver for the same dynamic graph $G$, supporting updates to $G$ in $O(U(n)+n)$ time. The data structure supports queries which for any given $b\in\mathbb{R}^{V}$ and $\epsilon>0$ in $\widetilde{O}(Q(n)\log 1/\epsilon)$ time return $x\in\mathbb{R}^{V}$ with $\|x-x^{*}\|_{\mathbf{L}_{G}}\leq\epsilon\|x^{*}\|_{\mathbf{L}_{G}}$ where $x^{*}$ is the exact solution for $\mathbf{L}_{G}x=b$.
Via Theorem 1.2 we have $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$, thus obtaining Theorem 1.4.
Proof B.5.
We run two separate data structures. The first is a dynamic spectral sparsifier Theorem B.2 which maintains a spectral sparsifier $H\approx\mathbf{L}_{G}$ with error $\epsilon=1/2$. We also run the dynamic OMv data structure for matrix $\mathbf{A}$ (adjacency matrix).
The update time is thus $\widetilde{O}(n+U(n))$ as the spectral sparsifier needs polylog time per edge updates, thus $\widetilde{O}(n)$ time for node updates.
Queries
When given a vector $b\in\mathbb{R}^{V}$ and error parameter $\epsilon>0$, we use $\mathbf{H}^{-1}$ as preconditioner for the linear system $\mathbf{L}_{G}x=b$. We can multiply with $\mathbf{H}^{-1}$ by running a static Laplacian system solver Lemma B.3 which takes $\widetilde{O}(n)$ time per product as $\mathbf{H}$ has $\widetilde{O}(n)$ edges. In particular, we let
$x^{0}=\mathbf{H}^{-1}b$
and then perform iterative refinement (Richardson iteration)
$x^{t+1}\leftarrow x^{t}-\mathbf{H}^{-1}(\mathbf{L}_{G}x^{t}-b).$
After $\widetilde{O}(\log 1/\epsilon)$ iterations we have $\|x^{t}-x^{*}\|_{\mathbf{L}_{G}}\leq\epsilon\|x^{*}\|_{\mathbf{L}_{G}}$.
Each iteration takes $\widetilde{O}(n+Q(n))$ time, where $\widetilde{O}(n)$ is the time for multiplying by $\mathbf{H}^{-1}$, and $Q(n)$ the time for multiplying by $\mathbf{L}_{G}$. Observe that $\mathbf{L}_{G}v=\mathbf{D}v-\mathbf{A}v$ where $\mathbf{D}$ is a diagonal matrix and $\mathbf{A}$ is an adjacency matrix. So the first product takes $O(n)$ time and the second product is handled by the dynamic OMv data structure in $O(Q(n))$ time. Together, this proves the lemma.
B.2 Effective Resistance
The following Theorem 1.5 is a corollary of the dynamic Laplacian solver from Theorem 1.4.
See 1.5
Proof B.6.
The effective resistance of a pair $u,v\in V$ is the energy of an electric flow, routing one unit of flow between $u$ and $v$. The energy of an electric flow $f\in\mathbb{R}^{E}$ is $\sum_{e}f_{e}^{2}$ and the electric flow is given by $f=\mathbf{B}\mathbf{L}^{\dagger}(e_{u}-e_{v})$ where $\mathbf{B}\in\{-1,0,1\}^{E\times V}$ is the incidence matrix of $G$ with arbitrary directions. Thus the effective resistance is
$\|\mathbf{B}\mathbf{L}^{\dagger}(e_{u}-e_{v})\|_{2}^{2}$ $=(e_{u}-e_{v})^{\top}\mathbf{L}^{\dagger}\mathbf{B}^{\top}\mathbf{B}\mathbf{L}^{\dagger}(e_{u}-e_{v})$
$=(e_{u}-e_{v})^{\top}\mathbf{L}^{\dagger}(e_{u}-e_{v}).$
This can be computed via the dynamic Laplacian system solver in Theorem 1.4.
B.3 Triangle Detection
See 1.6
The result follows form the following lemma for $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$ from Theorem 1.2.
Lemma B.7.
Let $G$ be a dynamic graph and let $\mathbf{A}$ be its dynamic adjacency matrix. Assume there is a dynamic OMv data structure for this $\mathbf{A}$ with update time $U(n)$ and query time $Q(n)$.
Then there exists a dynamic triangle detection algorithm for the same dynamic graph $G$, supporting updates in $O(U(n)+Q(n))$ time.
Proof B.8.
Let $\mathbf{A}$ be the adjacency matrix of the graph, then the diagonal entries $(\mathbf{A}^{3})_{v,v}$ are non-zero if and only if $v$ participates in a triangle. Thus by maintaining the sum $\sum_{i}(\mathbf{A}^{3})_{i}$ we can detect if there is any triangle in the graph.
We maintain the sum as follows: Let $\mathbf{A}$ be the incidence matrix before, and $\mathbf{A}$ after a node $v$ is updated.
Then
$\sum_{i}(\mathbf{A}^{\prime 3})_{i}=(\sum_{i}(\mathbf{A}^{3})_{i})+3(\underbrace{(\mathbf{A}^{\prime 3})_{v,v}-(\mathbf{A}^{3})_{v,v}}_{\text{difference in \# of triangles $v$ participates in}}).$
This is because each triangle $v$ participates in is counted thrice in the sum (once for each vertex of the triangle). Thus we can maintain the sum by computing $(\mathbf{A}^{\prime 3})_{v,v}$ and $(\mathbf{A}^{3})_{v,v}$. Observe that this is the vector-matrix-vector product of (i) the $v$th row of $\mathbf{A}$, (ii) matrix $\mathbf{A}$, (iii) and $v$-th column vector of $\mathbf{A}$. Thus it can be computed in $O(Q(n))$ time, with an extra $U(n)$ time to update $\mathbf{A}$ to $\mathbf{A}^{\prime}$.
B.4 Single-Source Shortest Paths and k-Center
Lemma B.9 (van den Brand et al. (2022)).
Assume there is a dynamic distance oracle for unweighted undirected graphs with the following operations.
An initialization procedure that is given $G=(V,E)$ and a threshold $1\leq d\leq n$.
A node update operation with update time $U(n,d)$.
A query operation which for any source node, given during the query, returns the $d$-bounded single source distances in $O(Q(n,d))$. That is, return for each $v\in V$ the distance if it is at most $d$, or $\infty$ if the distance is $>d$.
Then for any $1\leq k\leq n$ and $\epsilon>0$ there exists a dynamic $(1+\epsilon)$-approximate single source distance data structure on unweighted undirected graphs. It supports node updates in
$O\left(U\left(n,\frac{4}{\epsilon}\right)+\frac{k}{\epsilon}\cdot Q\left(n,\frac{4}{\epsilon}\right)+\frac{n^{2}}{k}\right)$
time.
See 1.7
Proof B.10.
We first describe how to obtain a dynamic distance oracle that yields distances up to $1/\epsilon$. This is done by running the dynamic OMv data structure of Theorem 1.2 on the adjacency matrix of $G$. Then single source distances up to $1/\epsilon$ from source vertex $u$ can be computed by repeatedly multiplying $w\leftarrow\mathbf{A}w$ for initial $w=e_{u}$, i.e., we compute $\mathbf{A}^{i}e_{u}$ for $i=1,2,...,1/\epsilon$. The smallest $i$ with $(\mathbf{A}^{i}e_{u})_{v}\neq 0$ is the smallest number of steps to reach from $u$ to $v$, i.e., the distance between the vertices.
This data structure internally constructs the a $\Delta$-labeled spanning tree of the boolean matrix $\mathbf{A}$, where the weight of the tree gives us the time complexity per query.
Hence, while we do not know the corrupted VC-dimension $d$, we do know how much time each query is going to take.
So when running Lemma B.9 on the above distance data structure, we can pick $k=n/\sqrt{T}$ where $T\geq n$ is the weight of the spanning tree. Thus, we get the update time
$\widetilde{O}\left(n+\frac{k}{\epsilon}T+nk+\frac{n^{2}}{k}\right)$ $=\widetilde{O}\left(n+\frac{n\sqrt{T}}{\epsilon}+\frac{n}{\sqrt{T}}\right)$
$=\widetilde{O}(n^{2-\frac{1}{2d}}),$
which proves the claim.
See 1.8
This result is a corollary of Theorem 1.7, as $k$-center can be reduced to computing $k$-single source distances.
Lemma B.11 (Theorem 7.3 by Cruciani et al. (2024)).
Given a graph $G=(V,E)$, a positive parameter $\epsilon\leq 1/2$, and a fully-dynamic data structure that maintains $(1+\epsilon)$-approximate single source distances with update time $T(n,m,\epsilon)$ and $Q(n,m,\epsilon)$ query time, there is a dynamic algorithm that maintains a $2(1+4\epsilon)$-approximate solution to fully-dynamic $k$-center in time $O(T(n,m,\epsilon)+k\cdot(Q(n,m,\epsilon)+n))$.
Appendix C Characterization of matrices with constant VC-dimension
In this section, we provide a characterization of matrices with constant VC-dimension. Particularly, we prove section 1 and state a number of examples of other constant VC-dimension matrices.
Fact 1.
Consider a hereditary class of $0/1$-matrices $\mathcal{M}$, meaning that $\mathcal{M}$ is closed under row/column deletion (i.e., each matrix $\mathbf{M}\in\mathcal{M}$ is still in $\mathcal{M}$ after deleting a row or column). If $\mathcal{M}$ is non-trivial, (i.e., does not contain every possible matrix), then there exists an absolute constant $c\in\mathbb{N}$ such that the VC-dimension of any matrix in $\mathcal{M}$ is at most $c$.
Fact 1 provides a structural characterization of some matrices with constant VC-dimension, and it is an analogous result to the fact that non-trivial hereditary classes of graphs have low VC-dimension. To prove Fact 1, we introduce the notion of a bipartisation of a graph.
Definition C.1.
A bipartite graph $H=(A\cup B,E)$ is a bipartisation of graph $G$ if removing all edges in $A$ and $B$ in $G$ yields $H$ for some partition $A,B$ of $V(G)$. A graph class $\mathcal{G}$ is said to be hereditary if it is closed under induced subgraphs.
Lemma C.2 (Lemma 3.4 of Bousquet et al. (2015)).
For any hereditary class $\mathcal{C}$ of graphs and for any bipartite graph $H$, if the graphs in $\mathcal{C}$ have infinite VC-dimension, then $\mathcal{C}$ contains a graph $G$ whose bipartisation is $H$.
By transposition, we get the following corollary:
Corollary C.3.
The equivalent contrapositive statement is that for any hereditary class $\mathcal{C}$ of graphs and for any bipartite graph $H$, if $C$ does not contain a graph $G$ whose bipartisation is $H=(A\cup B,E)$ where $A\cup B=V(G)$, then the graphs in $C$ have finite $\mathrm{VC}$-dimension.
Lemma C.4.
The $\mathrm{VC}$-dimension of any matrix in $\mathcal{M}$ is upper bounded by a finite constant.
Proof C.5.
Since $\mathbf{M}\in\mathcal{M}$ is Boolean, we can treat it as a bipartite graph $G$, with rows representing left vertices and columns representing right vertices. In particular, we can treat $\mathcal{M}$ as a class of bipartite graphs closed under vertex deletion (row deletion = deleting a left vertex, column deletion = deleting a right vertex). Thus we can apply lemma C.2 to $\mathcal{M}$.
Since $\mathcal{M}$ is non-trivial, there must be some graph $H$ not contained in it. For this excluded graph $H$, define $H^{\prime}$ as the same graph but add one extra left vertex and one extra right vertex. Connect this new left vertex to every vertex on the right, and connect the new right vertex to every vertex on the left. Note that since $H$ was not contained in $\mathcal{M}$ and it is closed under vertex deletion, $H^{\prime}$ also cannot be contained in $\mathcal{M}$.
Now assume there is $G\in\mathcal{M}$ that contains $H^{\prime}$ as bipartization. Then there is a partition $A,B$ of $V$, such that keeping only the edges between $A$ and $B$ of $G$, results in $H^{\prime}$. Because of the left vertex in $H^{\prime}$ that connects to all right vertices, and the right vertex connecting to all left vertices, we must have that $A$ are all left and $B$ are all the right vertices of $G$ (or the other way around). But that means no edges were removed from $G$ when restricting to edges between $A$ and $B$, since a bipartite graph by definition only has edges connecting left and right vertices. So $G=H^{\prime}$ and $H^{\prime}\in\mathcal{M}$. This is a contradiction because by construction of $H^{\prime}$, it does not exist in $\mathcal{M}$.
We conclude, that set of bipartite graphs corresponding to $\mathcal{M}$ does not contain any graph whose bipartisation is $H^{\prime}$. Hence by Corollary C.3 the VC-dimension must be a finite constant. Note that this is equivalent to the VC-dimension of the adjacency matrices being bounded, but $\mathcal{M}$ are not adjacency matrices. The adjacency matrices are of form
$\begin{bmatrix}0&\mathbf{M}\\ \mathbf{M}^{\top}&0\end{bmatrix}$
for matrices $\mathbf{M}\in\mathcal{M}$. We already argued that each adjacency matrix has constant VC-dimension, and since removing rows does not increase the VC-dimension (it is equivalent to deleting sets), this implies $\mathbf{M}$ also has constant VC-dimension. In conclusion, each matrix in $\mathcal{M}$ has constant VC-dimension.
C.1 Examples of low VC-dimension matrices
We provide some broad families of matrices with constant VC dimension.
Adjacency matrices of $H$-minor free graphs.
Eppstein (1995) showed that $H$-minor free directed graphs (where the minor can be obtained through vertex deletion, edge deletion, and edge contraction) has $\mathrm{VC}$-dimension at most $|V(H)|-1$. As an immediate consequence, since bipartite-graphs are triangle ($K_{3}$) free, the VC-dimension of any bipartite graph is $2$. As a consequence, given any adjacency matrix $\mathbf{A}$ of a $H$-minor free graph $G$, for any $k\geq 1$, $\mathbf{A}^{k}$ is the adjacency matrix of a $H$-minor free graph $G^{\prime}$ (since $\mathbf{A}^{k}$ corresponds to the reachability graph where an edge $(u,v)\in E(G^{\prime})$ corresponds to the existence of a $k$-length walk between $u$ and $v$ in $G$). Therefore, the VC-dimension of $\mathbf{A}^{k}$ is also at most $|V(H)|-1$.
Adjacency matrices of interval graphs.
Interval graphs are the intersection graphs of a family of intervals on the real line. Then, since the hypothesis class of intervals on the real line =$\mathcal{H}=\{\mathbbm{1}_{a,b}:a\leq x\leq b\}$ has VC-dimension $2$ (any two points can be shattered by choosing an interval that includes one or both points, but no set of three points can always be shattered), every interval graph has $\mathrm{VC}$-dimension at most $2$.
Boolean kernel matrices.
A matrix where each column $i$ (and each row) receives some label $\ell_{i}\in L$ from some (possibly infinite) set of possible labels $L$. Then, suppose each entry of $\mathbf{M}$ satisfies $\mathbf{M}_{i,j}=f(\ell_{i},\ell_{j})$ for some function $f:L\to\{0,1\}$. For example, the labels could be points in $L=\mathbb{R}^{d}$ with $f$ being indicator that the distance is at most some threshold. If $f,L$ is fixed (i.e., independent of the number of rows/columns) then these graphs are closed under row/column deletion and thus have bounded VC-dimension by Section 1.
Shortest-path structures
Let $G$ be an undirected graph with non-negative edge weights and let $S$ be a subset of its shortest paths such that, for every pair $(u,v)$ of distinct vertices, $S$ contains exactly one shortest path between $u$ and $v$. Marie de Lima et al. (2023) defines a range space associated with $S$ and proves that its VC dimension is 2.
Adjacency matrices of planar graphs.
Let $G$ be a planar graph. Then, there is a planar drawing of $G$, and deleting vertices and edges retains this property. Therefore, these graphs are closed under row/column deletion and thus have bounded VC-dimension by Section 1.
Adjacency matrices of semi-algebraic graphs of bounded description complexity.
By the Milnor-Thom theorem in real algebraic geometry (Matoušek, 2002) and the Sauer-Shelah lemma (Assouad, 1983), semi-algebraic graphs of bounded description complexity have constant VC-dimensions. In turn, this extends to adjacency matrices of string graphs where every two curves intersect.
Practical matrices.
Coudert et al. (2024) computes the VC-dimension of families of graphs that arise in practice, from protein interaction networks to autonomous Internet systems to web graphs. Even for such graphs with millions of nodes, the VC-dimension of the graphs are typically between $3$ and $8$.